### v1

# Language Bottleneck Models: A Framework for Interpretable Knowledge Tracing and Beyond

Abstract
Accurately assessing student knowledge is critical for effective education, yet traditional Knowledge Tracing (KT) methods rely on opaque latent embeddings, limiting interpretability. Even LLM-based approaches generate direct predictions or summaries that may hallucinate without any accuracy guarantees. We recast KT as an inverse problem: learning the minimum natural-language summary that makes past answers explainable and future answers predictable. Our Language Bottleneck Model (LBM) consists of an encoder LLM that writes an interpretable knowledge summary and a frozen decoder LLM that must reconstruct and predict student responses using only that summary text. By constraining all predictive information to pass through a short natural-language bottleneck, LBMs ensure that the summary contains accurate information while remaining human-interpretable. Experiments on synthetic arithmetic benchmarks and the large-scale Eedi dataset show that LBMs rival the accuracy of state-of-the-art KT and direct LLM methods while requiring orders-of-magnitude fewer student trajectories. We demonstrate that training the encoder with group-relative policy optimization, using downstream decoding accuracy as a reward signal, effectively improves summary quality.
1 Introduction
The Challenge of Estimating a Student’s Knowledge State.
A fundamental objective in education is accurately assessing what a student knows, identifying misconceptions, and understanding how their knowledge evolves over time posner1982accommodation ; larkin2012misconceptions ; chen2020impact . Teachers intuitively achieve this through diagnostic reasoning: by observing students’ answers, they infer not merely correctness, but deeper patterns reflecting conceptual mastery or specific misunderstandings. For instance, a student consistently misapplying algebraic properties signals conceptual gaps requiring targeted intervention. Replicating this nuanced interpretability within AI would profoundly enhance personalized education at scale, augmenting human teaching by providing transparent, actionable insights into learners’ knowledge states (khosravi2022explainable, ).
Building algorithms that can learn the knowledge state of students as they interact with educational material is an active field of study known as Knowledge Tracing (KT) (corbett_knowledge_1994, ; shen2024survey, ) (see Appendix A for a detailed introduction to KT). However, accurately modeling a student’s knowledge state in a way that is both predictive and interpretable remains a significant open challenge. Traditional KT methods (piech_deep_2015, ; zhang2017dynamic, ; pandey2019self, ; ghosh2020context, ) rely primarily on statistical, embedding-based representations. Such models lack transparency, representing knowledge states as high-dimensional latent vectors that provide limited interpretability or actionable insights for teachers and learners (Figure 1B). A truly impactful AI assistant in education must go beyond opaque predictions, explicitly articulating what the student understands and what they do not, in an accurate and interpretable manner.
Limitations of Existing LLM-based approaches for KT.
Recent approaches leveraging large language models (LLMs) have shown that LLMs can produce sensible predictions of students’ future behavior when provided with relevant information (li2024explainable, ; kim2024beyond, ), help mitigate the KT’s cold-start problem (lee2024language, ) and be finetuned to improve accuracy on KT tasks (wang2025llm, ). Nonetheless, these LLM-based methods still lack rigorous interpretability, as they either treat the model as a black-box or rely on free-form explanations that are susceptible to hallucination (bender2021dangers, ). Without ensuring faithful and grounded summaries, LLM-generated explanations fall short as reliable educational tools, emphasizing the need for a structured and principled foundation for generating interpretable, reliable, and adaptive student knowledge representations.
Figure 1: Language Bottleneck Models for Knowledge Tracing. The observed questions and answer $\mathcal{D}=\{\mathcal{X},\mathcal{Y}\}$ are caused by a certain knowledge state $\mathcal{S}$ held by the student when answering questions. The language bottleneck aims to approximate this knowledge state in natural language ($\tilde{\mathcal{S}}$). Then, the predicted knowledge state is used to predict both the observed historical data $\mathcal{X}$ and future questions $\mathcal{Y}$.
Knowledge Tracing as an Inverse Problem: Towards Concise and Interpretable Summaries.
Existing Knowledge Tracing methods typically approach student modeling as a forward prediction task, using latent statistical embeddings to predict future responses. However, this approach provides limited insight into the underlying student knowledge. We reframe Knowledge Tracing as an inverse problem: viewing student responses as outcomes of an underlying knowledge state $\mathcal{S}$, manifested through the student’s cognitive process (Figure 1A). The goal is to recover an interpretable representation $\tilde{\mathcal{S}}$ of the student knowledge state. Crucially, answers to downstream questions are expected to arise from the same latent knowledge state—an assumption we discuss further in Section 2.2. A useful representation must therefore reconstruct observed answers, enable prediction of future responses, and remain interpretable and steerable for teachers.
Introducing Language Bottleneck Models (LBMs).
Motivated by the inverse-problem framing described above, we introduce Language Bottleneck Models (LBMs). As illustrated in Figure 1C, LBMs uses two Large Language Models (LLMs) as student knowledge state extractor (encoder) and interpreter (decoder), generating succinct textual summaries meant to capture the student knowledge state. Specifically, LBMs map sequences of student interactions $\mathcal{X}$ to a concise, natural-language summary $\tilde{\mathcal{S}}$. This summary then serves as the sole basis for reconstructing the observed responses $\mathcal{X}$ and accurately predicting future responses $\mathcal{Y}$. By design, this ensures that the generated explanations encompass precisely the information necessary to interpret past behavior and reliably forecast future student performance.
Contributions and Paper Structure.
We make three key contributions:
  • We reframe knowledge tracing as an inverse problem and formalize a minimal-description objective that yields accurate and interpretable summaries of student knowledge states.
  • We introduce Language Bottleneck Models (LBMs)—an encoder-decoder architecture that forces all predictive information through a concise natural-language summary.
  • We empirically validate LBMs on synthetic and real-world datasets against strong KT and direct-prompting baselines, and show how reinforcement finetuning can train the encoder using downstream prediction accuracy.
The remainder of this paper is organized as follows: Section 2 introduces the knowledge tracing problem and reframes it as an inverse problem. Section 3 presents our Language Bottleneck Model (LBM) framework, along with results motivating this approach, and training methods. Section 4 discusses related literature on knowledge tracing and interpretability. Section 5 presents experimental results validating the effectiveness of LBMs. Finally, Section 6 discusses the applicability of LBMs beyond knowledge tracing and concludes with directions for future research.
2 Knowledge Tracing as an Inverse Problem
In this section, we formally introduce the knowledge tracing problem. We then reframe it as an inverse problem—inferring interpretable knowledge states from observed responses, outlining key desiderata for such representations and the conditions under which this formulation is appropriate.
2.1 Knowledge Tracing: Formal Problem Definition
Knowledge Tracing (KT) is fundamentally a sequence modeling problem where the objective is to infer a student’s latent knowledge state based on observed interactions. Formally, let $\mathcal{X}=\{(q_{1},a_{1}),(q_{2},a_{2}),\ldots,(q_{t},a_{t})\}$ denote a sequence of question-answer pairs, where $q_{i}\in\mathcal{Q}$ represents a question from the question space $\mathcal{Q}$, and $a_{i}\in\mathcal{A}$ represents the student’s answer from the answer space $\mathcal{A}$. The goal of KT is to learn a function that predicts the probability of correctly answering a future question $q_{t+1}$ given the historical interaction sequence: $P(a_{t+1}|q_{t+1},\mathcal{X})$.
While this formulation captures the predictive aspect of KT, it fails to address a crucial dimension: the interpretability. An ideal KT system should not only predict accurately but also provide educators with an explicit, human-understandable representation of the student’s knowledge state—such as mastered concepts, knowledge gaps and misconceptions. The importance of interpretability for KT is further discussed in the Appendix.
2.2 Knowledge Tracing as an Inverse Problem: From Black–Box Predictions to Interpretable Summaries
The conventional KT formulation (§2.1) focuses on predictive accuracy but offers limited insight into the underlying knowledge that drives student behavior. We instead frame KT as an inverse problem: in many practical scenarios—such as a quiz completed in a single sitting—a student’s responses can largely be explained by a fixed underlying knowledge state, expressed through the student’s cognitive process of answering questions. Most observed behavior reflects this latent state, with occasional deviations due to noise—such as slips (careless mistakes) or guesses (lucky correct answers), as formalized in Bayesian KT corbett_knowledge_1994 . The goal, then, is to recover an interpretable representation $\tilde{S}$ of the knowledge state that accounts for the observed interactions and supports accurate future predictions.
Framing KT in this way highlights three desiderata. (1) Self-consistency: the inferred summary $\tilde{S}$ should reconstruct the past by making the observed answers in $\mathbf{X}$ likely under a decoder. (2) Predictive sufficiency: $\tilde{S}$ must retain all information necessary to accurately predict future responses to unseen questions. (3) Human interpretability and steerability: the summary should be readable and editable by domain experts, enabling the diagnosis of student misconceptions and manual refinement or augmentation of $\tilde{S}$ to guide future predictions.
When is the inverse formulation appropriate? Although real learning is dynamic, many high–stakes diagnostic situations (end–of–unit quizzes, placement tests, or short tutoring sessions) span minutes rather than weeks. In such windows, it is reasonable to assume a quasi–static knowledge state. In this work, we explicitly focus on scenarios characterized by such quasi–static knowledge states. For longer horizons, the knowledge state might be better modeled as piecewise-static or smoothly evolving $S(t)$. We leave the extension of our framework to dynamically evolving knowledge states for future work.
2.3 Natural Language as the Interface
While a student’s knowledge state is inherently latent, natural language offers a flexible and interpretable medium for expressing it. Unlike fixed concept vocabularies, natural language accommodates open-ended reasoning patterns, misconceptions, and levels of mastery. In this work, we propose to use concise natural language summaries as proxies for student knowledge, designed to capture the latent cognitive state in a way that is both human-readable and predictive. The next section introduces a model family that operationalizes this idea.
3 Language Bottleneck Models
Figure 2: Zero-shot knowledge state encoding compared to ground truth. LLM models with different capabilities are prompted to write a summary of the knowledge state (top left panel) of a student, given 100 observed questions and answers provided in text. The ground truth knowledge state (top right panel) describing the student behavior has three main components: constructs mastered, constructs not mastered, and misconceptions. All three models capture the construct mastery correctly, but are not able to capture any misconception, beside o1 which notices the negative numbers misconception (bottom row). Bottom right symbols correspond to components in Figure 1.
3.1 Formal Definition
To tackle the inverse problem formulation of KT introduced in Section 2, we propose Language Bottleneck Models (LBMs): a two-stage architecture for inferring interpretable and accurate knowledge states from a student’s past interactions.
Let $X_{s}\subseteq X$ denote a subset of observed student responses, and let $Y_{s}\subseteq Y$ denote a set of future questions whose answers the model aims to predict. An LBM is defined by:
  • An encoder $f_{\theta}:X_{s}\rightarrow S$, which maps a history of interactions to a textual summary $S$ of the student’s implicit knowledge state.
  • A decoder $g:S\rightarrow\hat{Y}_{s}$, which uses only the textual summary to predict the student’s answers to new questions.
The explicit use of natural language as a bottleneck encourages the model to produce summaries that balance brevity with completeness. If the language summary omits essential details, predictions of future responses will suffer; conversely, if the summary is overly detailed or redundant, its utility for educators decreases. Thus, this design ensures the resulting textual summary is simultaneously concise, informative, and directly interpretable by humans.
The following sections motivate this approach and outline how LBMs can be trained effectively.
3.2 Motivating Observations
Before detailing our training procedure, we highlight two motivating observations that justify the LBM framework and its constraints.
Observation 1: Summarizing a knowledge state is challenging. Figure 2 compares summaries produced by different LLMs when prompted to describe a student’s knowledge state from 100 question-answer pairs from a synthetic dataset (see Section 5 for details). While all models capture high-level skill mastery, only one correctly identifies a misconception (errors with negative numbers) out of the four existing ones, illustrating that standard summarization approaches often miss crucial latent patterns.
Figure 3: Accuracy on synthetic dataset given ground truth knowledge state summaries.
Observation 2: Given a good summary, capable LLMs can decode with high fidelity. Figure 3 evaluates the performance of different decoder models when given access to the ground truth summary of the student’s latent knowledge state (an example of a ground-truth summary is shown in Figure 2). Stronger models like GPT-4o achieve nearly perfect accuracy ($98\%$), indicating that the bottleneck representation is indeed sufficient to drive effective downstream prediction—provided it encodes the right information.
Together, these observations motivate the inverse formulation: while generating summaries from raw interactions is hard, once obtained, these summaries are powerful inputs for reasoning and prediction. This validates our approach of learning encoders that optimize predictive utility via interpretable, compressed natural language representations.
3.3 Training Objective and Optimization
We train LBMs to produce informative and interpretable summaries of student knowledge states. As demonstrated above, strong decoders such as GPT-4o can achieve near-perfect accuracy when provided with accurate knowledge summaries. We therefore optimize the encoder $f_{\theta}$ to produce such summaries, while freezing the decoder $g$ to a strong LLM during training.
Summary generation and evaluation.
Given a student’s interaction history $X$, we sample a subset $X_{\text{enc}}\subseteq X$ of $N$ interactions to generate a summary $S=f_{\theta}(X_{\text{enc}})$. We evaluate this summary on two disjoint subsets: (i) $X_{s}\subseteq X_{\text{enc}}$: past questions for assessing reconstruction ability; and (ii) $Y_{s}\subseteq Y$: future questions for evaluating predictive accuracy. The summary $S$ is passed to the frozen decoder $g$, which outputs predictions $\hat{X}_{s}=g(S;X_{s})$ and $\hat{Y}_{s}=g(S;Y_{s})$.
Reward function.
The reward for a summary $S$ is defined using a function $\Phi$ that balances multiple criteria:
$\mathcal{R}(S;g)=\Phi\left(\texttt{Acc}(\hat{X}_{s},X_{s}),\;\texttt{Acc}(\hat{Y}_{s},Y_{s}),\;|S|,\;\Omega(S)\right),$ (1)
where $\texttt{Acc}(\cdot)$ measures accuracy, $|S|$ is the summary length, and $\Omega(S)$ enforces structural constraints (e.g., inclusion of a Misconceptions section). The function $\Phi$ balances these components using hyperparameters or indicator functions to enforce constraints as needed.
Optimization via GRPO.
To optimize $f_{\theta}$, we employ Group Relative Policy Optimization (GRPO) shao2024deepseekmath . For each input $X$, the encoder generates $G$ candidate summaries $\{S_{1},\dots,S_{G}\}$, each evaluated using $\mathcal{R}(S_{i};g)$. We compute group-relative advantage and update parameters:
$A(S_{i})=\frac{\mathcal{R}(S_{i};g)-\mu}{\sigma},\quad\nabla_{\theta}\mathbb{E}_{S\sim f_{\theta}(X)}[\mathcal{R}(S;g)]\approx\frac{1}{G}\sum_{i=1}^{G}A(S_{i})\cdot\nabla_{\theta}\log p_{\theta}(S_{i}|X),$ (2)
where $\mu$ and $\sigma$ are the mean and standard deviation of rewards within the group. This approach encourages the encoder to produce summaries that maximize the reward function defined in Eq. 1.
3.4 Steerability of the Estimated Knowledge States
A key advantage of LBMs is the ability for humans to steer the model’s behavior. We identify three mechanisms for human-model interaction (detailed in Appendix):
  • Prompt engineering the encoder. Directly shaping summary generation through input instructions.
  • Steering via reward signals. Incorporating human preferences during encoder training through the $\Omega(S)$ term in Eq. 1.
  • Augmenting with student-specific information. Educators can supplement the model with additional knowledge not present in observed data—either by augmenting encoder inputs or by editing the generated summary before decoding. This enables integration of recent classroom observations or specific misconceptions identified through in-person interactions.
4 Related Work
In this section, we review the literature related to our proposed approach, spanning traditional knowledge tracing, recent advances in LLM-based student modeling, and concept bottleneck models (CBMs), and text summarization models.
Traditional Knowledge Tracing
Traditional knowledge tracing (KT) methods have long been used to model student learning and predict future performance shen2024survey . Early works such as Bayesian Knowledge Tracing (BKT) corbett_knowledge_1994 model student mastery as a probabilistic process, while more recent methods like Deep Knowledge Tracing (DKT) piech_deep_2015 and Attentive Knowledge Tracing (AKT) ghosh2020context uses an attention-based architecture to learn contextual representation of questions and student knowledge states. Despite their strong predictive performance, these models represent a student’s knowledge state as an abstract, high-dimensional latent vector, which poses significant challenges in interpretability and actionable feedback for educators.
LLM-Based Knowledge Tracing
Recent studies have begun integrating Large Language Models (LLMs) into the KT framework. For example, li2024explainable demonstrated that LLMs are able to make sensible predictions about student responses when prompted with adequate information. Other works lee2024language ; kim2024beyond have studied how LLMs can help mitigate the cold-start problem compared to traditional KT approaches, while wang2025llm demonstrated state-of-the-art performances in KT by combining LLMs with sequence interaction models. However, these methods generally remain opaque: they either treat the LLM as a black-box, or rely on model-generated explanations that are susceptible to hallucination (bender2021dangers, ).
Concept Bottleneck Models
Concept Bottleneck Models (CBMs) koh2020concept improve interpretability through human-understandable concept activations as intermediates, with extensions exploring unsupervised concept learning oikarinen2023label , test-time interventions shin2023closer , and theoretical analyses of concept set design luyten2024theoretical . However, CBMs typically rely on finite predefined concept sets, limiting their applicability to complex tasks like knowledge tracing. Recently, yamaguchitoward introduced Explanation Bottleneck Models (XBMs), which use textual rationales as intermediates for vision classification, justifying a single known label per input. While our Language Bottleneck Models (LBMs) adopt this language bottleneck concept, they differ fundamentally. Unlike XBMs’ instance-specific, task-specific rationales, LBM summaries aim to capture an implicit knowledge states—as emphasized by our inverse problem formulation—and generalize to future, unknown questions. These requirements necessitate holistic, adaptable summaries rather than XBMs’ task-specific rationales, leading us to introduce Language Bottleneck Models (LBMs) as a distinct framework with broader applicability.
Contrast with Text Summarization Models
Recent text summarization models such as BART lewis2019bart , T5 raffel2020exploring , and PEGASUS zhang2020pegasus effectively produce concise summaries based on explicitly available textual content. However, these approaches differ fundamentally from knowledge tracing, where summaries must infer implicit student knowledge states not directly observable in the input. Standard summarization metrics (e.g., ROUGE, BLEU) rely on explicit reference summaries, making them unsuitable for evaluating latent knowledge inference tasks. In contrast, our Language Bottleneck Models generate textual summaries of the student’s implicit knowledge state, optimized for predictive accuracy on downstream questions rather than syntactic overlap with the observed input.
5 Experiments
We evaluate the effectiveness of our Language Bottleneck Models (LBM) in the context of Knowledge Tracing (KT), both with a controlled synthetic KT task involving basic arithmetic operations, and on real-world knowledge tracing datasets.
5.1 Experimental Setup
Datasets We evaluate on one synthetic and one real-world source. Synthetic arithmetic simulates learners answering addition, subtraction, multiplication, and division items; each student is assigned mastered skills, unmastered skills, and systematic misconceptions. The Eedi dataset wang2020diagnostic contains millions of question–answer pairs from UK secondary-school mathematics. More details about these datasets are provided in the Appendix.
Models The main experiments use GPT-4o and GPT-4o-mini, together with open-source Qwen 2.5 (3B, 7B) and Gemma 3 (12B, 30B) checkpoints. Unless noted, we run the instruct variants of each open-source model; hyper-parameters and prompt templates are provided in the Appendix.
Baselines We compare LBMs against (i) four standard knowledge-tracing methods—DKT piech_deep_2015 , DKVMN zhang2017dynamic , SAKT pandey2019self , and AKT ghosh2020context —implemented with the pyKT library liupykt2022 ; and (ii) direct prompting, where each LLM predicts answers from the full student history without a bottleneck. Training details for all baselines are given in Appendix.
5.2 Results
5.2.1 LBM vs Direct Prompting
Table 1: Accuracy results for Direct and LBM methods. Bold indicates LBM accuracy is no more than 2% below the Direct baseline.
synthetic eedi Direct LBM $\Delta$ Direct LBM $\Delta$ Qwen2.5-3B 0.525 0.627 0.102 0.347 0.372 0.025 Qwen2.5-7B 0.655 0.646 -0.009 0.647 0.583 -0.065 gemma-3-12b 0.790 0.791 0.001 0.580 0.625 0.045 gemma-3-27b 0.785 0.794 0.009 0.670 0.650 -0.020 gpt-4o-mini 0.795 0.782 -0.013 0.655 0.645 -0.010 gpt-4o 0.840 0.825 -0.015 0.703 0.682 -0.020
Figure 4: Evolution of accuracy with bottleneck length.
Table 1 compares the accuracy of LBM models to direct prompting variants across different LLMs. Despite the information bottleneck enforced by LBMs, they remain within $2\%$ of the accuracy obtained via direct prompting. Figure 4 demonstrates that this gap decreases for longer bottlenecks, highlighting a trade-off between conciseness and predictive accuracy.
Encoder/decoder variants
We evaluate LBMs with different models as encoders and decoders (results shown in the Appendix). Using the strongest model (gpt-4o) as encoder with other models from Table 1 as decoders yields accuracies $5-10\%$ higher than when the stronger model is used as the decoder. This confirms our hypothesis that extracting relevant information for the summary (encoding) is more challenging than predicting future answers given a summary (decoding).
5.2.2 LBM vs Knowledge Tracing Methods
Figure 5 compares the performance of LBM models to traditional Knowledge Tracing methods on the synthetic dataset. LBMs with strong LLMs achieve comparable accuracy to KT methods with significantly less training data. Since KT methods rely on statistical patterns, they require substantially more observations before reaching strong predictive power, while LBMs demonstrate effective generalization with minimal examples.
Figure 5: Accuracy of LBM vs Knowledge Tracing models on the Synthetic dataset. The x axis represents the total number of question seen by the model, ie the product of number of questions observed per student $T$ and the number of students available for training $N$ (note that N=1 for LBM methods). The y axis is the average accuracy 200 test student and 4 new questions to predict per student. The pre-trained models indicated for LBM are used as both Encoder and Decoder, with chain-of-thought prompting. Error bars report the standard error calculated across 10 runs for N=10 students and 3 runs otherwise.
5.2.3 Training LBM Encoders
As discussed in Section X, downstream accuracy serves as a reward signal to train encoders to produce increasingly accurate summaries. We demonstrate this by training a Gemma3-12B encoder with GRPO alongside a frozen Gemma3-27B decoder. Figure 6 shows the encoder progressively improving summary quality, quickly outperforming gpt-4o. The figure illustrates this progression through an example student who mastered all constructs except subtraction and fails any multiplication involving 9. The initial summary contains inaccuracies and misses this systematic misconception, while the final summary successfully captures the student’s complete knowledge state.
Figure 6: Training the encoder with the decoding accuracy on the synthetic dataset. Evolution of the test accuracy as a Gemma3-12B encoder is trained as described in Section 3.3, with a fixed Gemma3-27B decoder. Trained on $800$ training students and tested on $200$ students, with $|\mathcal{X}|=50$ questions per input trajectory and $|\mathcal{Y}|=20$ questions to predict per student. The bottom row show the evolution of the generated summary over the course of training for an example student. Text is colored green (exact), orange (approximate), or red (false) based on ground-truth.
Accuracy gains from training are stronger on harder cases
When stratifying students by difficulty (measured by number of misconceptions), we find training yields larger accuracy improvements for complex cases (see table in the Appendix). The model shows 10-11% increased accuracy relative to gpt-4o for students with 2 or 3+ misconceptions, compared to just 5-7% for students with 0-1 misconceptions. This suggests the encoder becomes particularly better at handling complex knowledge states.
5.2.4 Steering LBM behavior
We demonstrate multiple steering strategies described in Section 3.4 (details in Appendix). We find that providing explicit misconception information during encoder training produces significantly stronger learning effects than adding the same information at the decoder stage. This suggests the encoder leverages this additional context to better interpret patterns in student responses. We also show the encoder can be steered during training to explicitly mention misconceptions through reward signals, and illustrate how supplementing the bottleneck with information not present in the input can significantly improve decoder accuracy.
6 Discussion
Core contributions
Framing Knowledge Tracing as an inverse problem shifts focus from fitting statistical regularities to understanding the underlying reasons for student responses. Language Bottleneck Models operationalize this view by compressing interaction histories into explicit natural-language summaries that maintain high predictive power while being both interpretable and steerable by domain experts.
Why can’t we just prompt GPT-4o directly?
The split encoder-decoder architecture of LBMs offers three key advantages over direct LLM prompting: it creates a global student model with a single latent summary shared across all predictions rather than isolated per-question reasoning; it ensures faithful summaries through a closed-loop decoding objective that penalizes non-predictive summaries; and it provides an explicit interface layer that teachers can read, steer and intervene on.
Wider applicability
LBMs extend beyond education to any task requiring compact, human-readable summaries with predictive power. The minimal ingredients needed are: (1) a sequence of observations about an entity, (2) a need to predict future behaviors of that entity, and (3) value in having interpretable representations. For example, clinical decision support could distill patient data into textual state descriptions that forecast outcomes while remaining auditable; preventive maintenance could compress sensor logs into explanations predicting machine failure; customer success teams could summarize interaction histories to forecast churn.
Limitations
LBMs face several constraints: context length limitations imposed by underlying LLMs, which could be addressed through iterative summarization; the requirement for textual question content rather than just identifiers, limiting applicability to some KT datasets; and the substantial computational cost of running LLMs. We discuss limitations extensively in the Appendix.
Future work
Several promising extensions include: implementing iterative encoding to process longer inputs through sequential chunking; developing active sensing capabilities to select questions that maximize information about students’ knowledge states; adapting LBMs for evolving knowledge states to capture learning and forgetting over time; expanding beyond question-answer inputs to leverage richer data sources; and combining with pedagogical-alignment techniques like LearnLM (team2024learnlm, ). We discuss these points in more details in the Appendix. By bridging the gap between predictive power and human interpretability, Language Bottleneck Models represent a significant step toward AI systems that can genuinely augment human expertise in education and beyond.
Appendix A Introduction to Knowledge Tracing
In this section, we provide a more detailed introduction to Knowledge Tracing. For recent comprehensive reviews of the field, see [1, 25].
Knowledge Tracing (KT) is the task of modeling students’ knowledge states and predicting their performance on future questions based on their historical interactions with educational content. The goal is to dynamically track the acquisition, retention, and application of knowledge as students engage with learning materials.
A.1 Formal Problem Definition
Knowledge Tracing (KT) is fundamentally a sequence modeling problem where the objective is to infer a student’s latent knowledge state based on observed interactions. Formally, let $\mathcal{X}=\{(q_{1},a_{1}),(q_{2},a_{2}),\ldots,(q_{t},a_{t})\}$ denote a sequence of question-answer pairs, where $q_{i}\in\mathcal{Q}$ represents a question from the question space $\mathcal{Q}$, and $a_{i}\in\mathcal{A}$ represents the student’s answer from the answer space $\mathcal{A}$. The goal of KT is to learn a function that predicts the probability of correctly answering a future question $q_{t+1}$ given the historical interaction sequence: $P(a_{t+1}|q_{t+1},\mathcal{X})$.
In practice, each interaction $(q_{i},a_{i})$ in the dataset typically contains richer information. Questions $q_{i}$ often include metadata about the underlying knowledge constructs they assess (e.g., specific skills, concepts, or learning objectives). For example, a mathematics question might be tagged with constructs like "quadratic equations" or "algebraic factoring." Similarly, observed answers $a_{i}$ can vary in granularity—from binary correctness indicators to specific answer choices in multiple-choice questions, or complete textual/code responses for open-ended questions. Many KT datasets also capture additional contextual information such as response time or self-reported confidence scores, which can be leveraged by KT models.
The prediction target can also extend beyond binary correctness. Models may predict specific answer choices a student will select (see for example Task 2 of Eedi’s NeurIPS 2020 Education Challenge [30]), enabling more nuanced understanding of misconceptions. In domains like programming education, models might predict the exact code a student will write [14], capturing finer-grained knowledge states.
A.2 Knowledge Tracing Paradigms
Bayesian Knowledge Tracing (BKT).
Introduced by [5], BKT models student knowledge as a set of binary latent variables corresponding to specific skills or concepts. For each skill, a student is either in a "known" or "unknown" state. BKT employs a Hidden Markov Model to update these probabilities based on observed student responses, accounting for learning, guessing, and slipping in its probabilistic framework. This approach provides interpretable outputs, with each concept having a probability of mastery, but is limited in modeling the complex interdependencies between skills.
Deep Learning Approaches.
Deep Knowledge Tracing (DKT) [20] marked a paradigm shift by using recurrent neural networks to model student knowledge as continuous latent representations. Unlike BKT’s discrete skill probabilities, DKT captures complex patterns and skill relationships through high-dimensional embeddings, significantly improving prediction accuracy. Subsequent models have enhanced this approach, including: Dynamic Key-Value Memory Networks (DKVMN) [34] which incorporate memory components for better concept tracking; Self-Attentive Knowledge Tracing (SAKT) [19] and Attentive Knowledge Tracing (AKT) [6] which leverage attention mechanisms to capture long-range dependencies; and Graph-based Knowledge Tracing models [17, 33] which incorporate structural information into knowledge representations.
A.3 Current Challenges
Interpretability vs. Accuracy Trade-off.
Deep learning approaches typically achieve higher predictive accuracy but at the cost of interpretability. Their latent representations are opaque high-dimensional vectors that provide limited insight into specific concepts a student has mastered or misconceptions they hold. This limits the usability of these models in practice.
Why do we need interpretability?
At first glance, predictive accuracy alone might seem sufficient for KT—an accurate model could, for example, analyze student responses across many questions to identify overall strengths and weaknesses. However, even highly accurate models can yield misleading or fragile predictions if the underlying reasoning is opaque. For instance, [22] showed that predictive accuracy alone might obscure models relying on superficial patterns like repeated questions, limiting practical usability. Thus, without interpretability, accurate models may fail to meaningfully support educators in diagnosing and intervening in students’ learning processes.
Cold-start Problem.
KT systems struggle with new students or questions where historical data is limited or non-existent: for any new question added to the dataset, traditional models require the observation of answers from many students to learn an accurate latent representation for this question. This "cold-start" problem poses significant challenges for the practical deployment of KT systems, particularly in dynamic educational environments where new content and students are regularly introduced.
Appendix B Extended Results
B.1 Encoder/Decoder Variants (Section 5.2.1)
Table B1 shows the accuracy of different combinations of encoder-decoder models on the Synthetic dataset. The top row shows the result of using the strongest model evaluated (gpt-4o) as both encoder and decoder. Then, we vary either the encoder or decoder part of the LBM across other LLMs, and report the resulting accuracy. A clear pattern which emerges is that the resulting LBMs are stronger when gpt-4o is used as the encoder instead of the decoder. This implies that the task of accurately capturing knowledge state information in the bottleneck is harder than predicting answers to future questions when provided with a knowledge state summary.
Table B1: Performance of different LBM encoder-decoder combinations on the Synthetic dataset.
Encoder Decoder Accuracy
Strongest model gpt-4o gpt-4o 0.809
Strongest gpt-4o gpt-4o-mini 0.821
model gpt-4o google/gemma-3-12b-it 0.795
as gpt-4o Qwen/Qwen2.5-7B-Instruct 0.765
encoder gpt-4o Qwen/Qwen2.5-3B-Instruct 0.715
Strongest gpt-4o-mini gpt-4o 0.765
model google/gemma-3-12b-it gpt-4o 0.775
as Qwen/Qwen2.5-7B-Instruct gpt-4o 0.666
decoder Qwen/Qwen2.5-3B-Instruct gpt-4o 0.649
B.2 Traditional KT methods on the Eedi dataset
We run the four benchmark KT methods on the Eedi dataset with the same experimental setup as for Table 1 in the main paper (see Section D.2.2 in this appendix for dataset details). Table B2 shows the resulting accuracies, alongside the accuracy found with Direct and LBM models for different LLMs, reproduced from Table 1. The best KT model (AKT) slightly outperform the best LBM and Direct models with gpt-4o, while other KT models reach similar accuracies to LBM and Direct models with gpt models and Gemma3-27B.
Table B2: Average accuracy on the Eedi dataset for different model types. For each student, the accuracy is computed on the last four questions of the trajectory. KT models show mean $\pm$ SEM across 10 runs. Due to computational constraints, Direct and LBM results are reported from single runs.
Model Accuracy
KT Models
DKT $0.647\pm 0.005$
DKVMN $0.669\pm 0.004$
SAKT $0.680\pm 0.007$
AKT $0.720\pm 0.008$
Direct
Qwen2.5-3B 0.347
Qwen2.5-7B 0.647
Gemma3-12B 0.580
Gemma3-27B 0.670
GPT-4o-mini 0.655
GPT-4o 0.703
LBM
Qwen2.5-3B 0.372
Qwen2.5-7B 0.583
Gemma3-12B 0.625
Gemma3-27B 0.650
GPT-4o-mini 0.645
GPT-4o 0.682
B.3 Training LBM Encoders - Difficulty stratification (Section 5.2.3)
To investigate whether the accuracy gains seen during LBM training (cf Figure 6 in the main paper) vary across students in the dataset, we stratify students according to how many misconception they hold: 0, 1, 2 or 3+. Since each misconception represent additional "irregularities" in the student’s response pattern beyond simple mastery of constructs, this effectively stratifies different difficulty levels across students. Table B3 shows the change in accuracy relative to the gpt-4o baseline across difficulty levels.
Table B3: Relative difference in accuracy between trained Gemma3-12B and the GPT-4o baseline, with students grouped by number of misconceptions.
Accuracy (mean ± std) Relative Diff. (%)
# of Misconceptions Baseline Pre-training Post-training Pre Post $\Delta$
0   ($N=22$) 0.98 ± 0.05 0.92 ± 0.12 0.99 ± 0.05 -5.80 0.90 6.70
1   ($N=17$) 0.89 ± 0.09 0.87 ± 0.15 0.91 ± 0.09 -2.00 3.00 5.00
2   ($N=23$) 0.81 ± 0.16 0.80 ± 0.12 0.89 ± 0.08 -0.80 10.20 11.00
3   ($N=135$) 0.77 ± 0.12 0.74 ± 0.14 0.82 ± 0.11 -4.80 5.70 10.50
B.4 Steering LBM behavior
B.4.1 Steering via reward signals
To demonstrate the possibility to steer LBMs’ behavior via the reward signal, we consider an example use-case where a teacher would like the model to pay particular attention to potential misconceptions held by the student. Following the reward-shaping framework of Section 3.3, we augment the training objective with an additional term that explicitly encourages the model to surface misconceptions. Concretely, we set $\Omega(S)=\mathbbm{1}\!\bigl{[}\text{``misconception''}\in\tilde{\mathcal{S}}\bigr{]}$, where $\tilde{\mathcal{S}}$ is the textual bottleneck emitted by the LBM for student state $S$. This binary reward is added to the accuracy term and optimized with GRPO. Figure B1 confirms that the policy quickly internalizes this incentive: after only a handful of training steps, the proportion of bottlenecks that explicitly mention a misconception goes from less than 80% to >95%, demonstrating that the reward function provides an effective lever for shaping higher-level pedagogical behavior.
Figure B1: Presence of the word "misconception" in the LBM’s bottlenecks during training of the encoder model steered towards mentioning student misconceptions via the reward signal.
B.4.2 Augmenting LBMs with student-specific information to complement input data
To demonstrate how an LBM can be actively steered with additional information we run the following ablation on the Synthetic dataset: 1. Each student trajectory originally probes four constructs. For a given trajectory, we remove every question linked to one construct leaving the input data intentionally incomplete. 2. We then run the input data through the encoder and inject a single teacher-supplied sentence describing the student’s mastery of the missing construct directly into the model’s bottleneck representation (e.g., "The student has mastered addition except in the event of misconceptions.).
We repeat this procedure four times—once for each construct—running the LBM both with and without the additional sentence, and report the results Table B4. Naturally, the models complemented with the additional information of the student’s mastery of the construct missing in the input data outperform models provided only with the incomplete data. This small experiment illustrates a key advantage of LBMs: because the LBM compresses evidence into a text-based bottleneck, it can be complemented with additional information absent from the input data.
Table B4: Accuracy of a gpt-4o-based LBM on the Synthetic dataset while removing one construct from the input data, with and without providing information about the missing construct in the bottleneck. Each run is repeated across all four constructs; mean and standard deviation are reported.
Without additional bottleneck information With additional bottleneck information
0.749 ± 0.007 0.791 ± 0.013
B.4.3 Augmenting with student-specific information to assist LBM training
Finally, we demonstrate how providing additional information to the LBM can assist with the training process. We train two identical LBMs while providing 2 misconceptions either to the encoder as part of the input data, or to the decoder as part of the bottleneck. The models are trained for one epoch on $800$ students. We then evaluate the resulting models without any additional information provided. Table B5 shows the accuracy before/after training for both models. The model where additional information was provided to the encoder reaches $84\%$ accuracy after one training compared to $80\%$ when it is provided to the decoder. This suggest that the additional information facilitates training of the encoder.
Table B5: Relative difference in accuracy between trained Gemma3-12B and the gpt-4o baseline, with students grouped by number of misconceptions.
Before training After 1 epoch training
Information in X 0.802 ± 0.001 0.840 ± 0.004
Information in S 0.794 ± 0.002 0.799 ± 0.006
B.5 Example bottlenecks on the Eedi dataset
Figure B2 shows example bottlenecks produced by a gpt-4o-based LBM for a student in the Eedi dataset. The input data is composed of 30 questions from various constructs, and the decoder predicts 4 test questions. As shown Figure 4 in the main paper, a shorter bottleneck of 128 tokens constrains the expressive power of the model, and in this example the resulting predictions fail on two of the test questions. A longer bottleneck of 256 or 512 tokens allows for more nuance and details in the bottleneck, leading to the decoder correctly predicting all four test questions for this student.
Figure B2: Example bottlenecks produced by an LBM with gpt-4o as both encoder and decoder for a student in the Eedi dataset, with varying bottleneck lengths. The input data provided to the encoder is composed of 30 questions across various constructs, and the decoder predicts 4 test questions using the bottleneck.
Appendix C Extended Discussion
C.1 Extended Discussion
Knowledge Tracing as an Inverse Problem
The reframing of Knowledge Tracing as an inverse problem represents a fundamental shift in how we conceptualize student modeling. Traditional KT methods adopt a forward modeling approach, where the goal is to learn a mapping from observed interactions to future performance predictions. While effective for prediction, this treats the student’s knowledge state as an opaque latent variable that serves primarily as an intermediate computational step. In contrast, our inverse problem formulation explicitly seeks to recover the underlying knowledge state that generates observed behaviors, making interpretability the primary objective rather than an afterthought. This aligns with educational assessment principles, where the main objective is to understand what a student knows, rather than merely predicting future behavior.
Parallel with Kolmogorov Complexity
Our inverse problem formulation of Knowledge Tracing draws a natural parallel with algorithmic information theory. The Kolmogorov complexity $K(x)$ of a string $x$ is defined as the length of the shortest program that outputs $x$ when run on a universal Turing machine. Traditional KT methods focus on learning statistical regularities between questions and responses, without looking for parsimonious explanations for observed behavior. In contrast, LBMs explicitly search for minimal natural language descriptions that can both reconstruct past interactions and predict future responses. The knowledge state summary $\mathcal{S}$ is akin to the minimal program, and the decoder LLM functions is akin to the universal interpreter that unpacks $\mathcal{S}$ to generate the observed question-answer patterns. This connection suggests that effective knowledge state summaries should capture the algorithmic essence of student behavior—the underlying "program" of knowledge and misconceptions that generates observable responses—rather than merely fitting surface-level patterns. While true Kolmogorov complexity is uncomputable, LBMs approximate this ideal through the natural language bottleneck constraint, encouraging summaries that balance compression with predictive power.
Computational Architecture and Design Choices
The encoder-decoder architecture reflects a deliberate separation of concerns: extraction versus interpretation of knowledge states. Our experiments demonstrate that these two tasks have different intrinsic complexity, with encoding (summary generation) being more challenging than decoding (prediction from summaries). This finding has practical implications for model selection and computational resource allocation. The use of GRPO for encoder training represents a novel application of reinforcement learning to interpretable AI, where the reward signal directly measures the downstream utility of explanations rather than their surface-level quality.
C.2 Extended Limitations
Context Length Constraints
A practical limitation of LBMs is the context length restrictions of current LLMs. Eventhough modern models can handle thousands of tokens, comprehensive student histories in real educational settings can easily exceed these limits. As student trajectories grow more complex and include more information, the context required for the encoder might go beyond what current LLMs are capable of. A natural solution could be an iterative encoding process, where the encoder iterates over the text bottleneck while going over the input data by windows.
Computational Cost and Resource Requirements
Since LBMs involve two LLMs as the encoder and decoder, they are typically more computationally intensive to run than traditional KT methods. Each inference call with LBMs requires two LLM calls (encoding and decoding) with extensive context windows. Training LBMs with GRPO is particularly costly, as it requires iteratively generating and evaluating multiple candidate summaries per training example. This cost structure may limit the practical deployment of LBMs to high-stakes educational contexts where the interpretability benefits justify the computational expense.
Dependence on Textual Question Content
Many existing KT datasets provide only question identifiers rather than full question text. This limits the direct applicability of LBMs, as they require the questions content in order to generate meaningful summaries about student knowledge. While question content is increasingly available in modern educational platforms, this dependency creates a barrier to applying LBMs to historical datasets or systems that rely primarily on item response theory frameworks.
Quasi-Knowledge State Assumption
The quasi-static knowledge state assumption underlying our inverse problem formulation, while reasonable for short diagnostic sessions, does not cover longer time horizons. Real learning involves continuous knowledge acquisition, forgetting, and misconception evolution that our current framework cannot capture. This limitation restricts current LBMs to diagnostic rather than contexts of assessment throughout the learning process. The assumption also fails to account for contextual factors (fatigue, motivation, external stressors) that can significantly impact student performance within even short sessions.
C.3 Extended Future Work
Iterative Encoding Strategies
A natural extension to address context length limitations involves iteratively applying the encoder while chunking input data. At each step, the encoder would process a portion of the interaction history $X$ along with one or more previously generated bottleneck summaries, creating an updated summary that incorporates new information from the latest chunk. This approach could maintain both detailed recent context and compressed historical patterns, enabling LBMs to handle arbitrarily long student trajectories through sequential refinement of knowledge state representations.
Active Learning and Question Selection
Building on iterative encoding, active learning strategies could provide encoders with the most informative input questions. Rather than processing all available interactions, the system could strategically select questions that maximize information gain about uncertain aspects of student knowledge. This could leverage uncertainty quantification methods or use the encoder LLM itself to recommend the most diagnostically valuable questions. Such active sensing would improve both efficiency and diagnostic power by focusing computational resources on the most insightful student responses.
Dynamic Knowledge State Modeling
Extending LBMs to handle evolving knowledge states represents a significant next step necessary to account for progressive learning and forgetting effects. A natural relaxation of the static knowledge state assumption involves "piecewise-constant" knowledge states that can evolve between question sessions but remain static within sessions. This extension poses exciting challenges in developing training objectives that balance within-session consistency with between-session learning dynamics, potentially requiring new approaches to temporal modeling in natural language representations.
Multimodal Input Integration
Question-answer pairs represent a relatively narrow information source for inferring student knowledge states. Richer data sources such as student-tutor interactions, self-reported explanations for answers, timing patterns, or hint usage could provide deeper insights into student understanding. While LBMs naturally extend to text-based inputs, future work should investigate optimal strategies for combining these varied data sources and evaluate their relative contributions to knowledge state inference accuracy and interpretability.
Pedagogical Alignment
Recent work like LearnLM [28] demonstrates the potential for making LLMs more pedagogically aligned. Incorporating similar pedagogical principles into LBM components could help encoders better interpret educational interactions and generate summaries that align with expert teaching practices. This might involve training on educator-annotated examples, incorporating educational taxonomies into summary structure, or using reward functions that emphasize pedagogically relevant aspects of student knowledge states. Such alignment could bridge the gap between computational convenience and educational validity.
Appendix D Experimental Details
D.1 Models
LLM models
We evaluate the following closed- and open-source models:
  • GPT-4o and GPT-4o-mini[2];
  • Qwen2.5-3B and Qwen2.5-7B[29];
  • Gemma3-12B and Gemma3-27B[27].
For open-source models we extract the activation logits of the “Yes” and “No” tokens and return the higher value. For closed-source models we prompt for a “Yes” or “No” answer and parse the text output.
Knowledge Tracing models
We evaluate the following four Knolwedge Tracing models:
  • DKT [20]
  • DKVMN [34]
  • AKT [6]
  • SAKT [19]
For each one of these models, we use the pyKT implementation [15] with default hyperparameters.
D.2 Datasets
D.2.1 Synthetic
Our synthetic dataset simulates students answering basic arithmetic problems. Each student is characterised by (i) mastered skills, (ii) unmastered skills, and (iii) systematic misconceptions. Questions are arithmetic operations—addition, subtraction, multiplication, or division (rounded to the nearest integer)—between two operands drawn from $[0,15]$ for addition/subtraction or $[1,10]$ for multiplication/division.
Misconception pool.
For every student we sample misconceptions uniformly at random from:
  • forgets to carry in addition;
  • fails multiplications involving the number $x$ with $x\sim\mathcal{U}(6,9)$;
  • fails any operation involving the number $x$ with $x\sim\mathcal{U}(6,9)$;
  • fails whenever an operand $>10$;
  • always rounds division results down;
  • fails with negative numbers.
Generation parameters.
  • Number of students: $2000$;
  • Number of questions: $5000$;
  • Questions answered per student: $210$.
Figure D1 shows the histogram of the number of misconceptions per student across the $2000$ students of the dataset.
Figure D1: Distribution of the number of misconceptions per student in the Synthetic dataset.
D.2.2 Eedi
The Eedi dataset is analogous to the one publicly shared via the NeurIPS 2020 Education Challenge [30] organised by Eedi, but additionally includes the text of each question.
Although many trajectories in the Eedi dataset span several days or months, we retain only single-session sequences to satisfy the quasi-static-knowledge assumption:
  • minimum response time: $5\,\mathrm{s}$;
  • maximum gap between answers: $3\,\mathrm{min}$;
  • minimum trajectory length: $40$ questions.
This filtering yields a total of $623$ individual trajectories with 40+ questions.
D.3 Results
D.3.1 Table 1
Synthetic dataset
  • $|\mathcal{X}|=50$
  • $|\mathcal{Y}|=4$
  • Test students: $200$
  • Bottleneck size: $128$ tokens
  • Chain-of-thought prompting: No
Eedi dataset
  • $|\mathcal{X}|=30$
  • $|\mathcal{Y}|=4$
  • Test students: $100$
  • Bottleneck size: $512$ tokens
  • Chain-of-thought prompting: No
D.3.2 Figure 4
Same set-up as Table 1; only the bottleneck size varies (see $x$-axis).
D.3.3 Figure 5
Latent Bottleneck Models (LBMs)
  • $|\mathcal{X}|=50$, $|\mathcal{Y}|=4$
  • Test students: $200$
  • Bottleneck size: $256$ tokens
  • Chain-of-thought prompting: Yes
  • Encoder & decoder: identical (varies by row in the table)
Knowledge-tracing baselines DKT [20], DKVMN [34], AKT [6], and SAKT [19] are evaluated using the pyKT implementation [15], modified to compute accuracy on $|\mathcal{Y}|=4$ questions, in order to be comparable with our LBM models. For each method we keep the pyKT default hyperparameters.
D.3.4 Figure 6
  • Dataset: Synthetic
  • $|\mathcal{X}|=50$, $|\mathcal{Y}|=20$
  • Train / test students: $800/200$
  • Bottleneck size: $128$ tokens
  • Chain-of-thought prompting: No
  • Encoder (trained): Gemma3-12B
  • Decoder (frozen): Gemma3-27B
Training hyper-parameters.
[…]
  • $G=5$
  • Learning $\text{Learning rate}=1\times 10^{-4}$
  • $\beta=0.04$
  • Optimiser: Adam (adamw_torch, default settings)
  • LoRA configuration
    • $r=16$, $\alpha=16$
    • target_modules = ["q_proj", "k_proj", "v_proj", "o_proj", "gate_proj", "up_proj", "down_proj"]
    • Dropout: $0.05$
D.3.5 Encoder–Decoder Variants
Identical to the Table 1 set-up, except for the choice of encoder and decoder LLMs.
D.3.6 Difficulty Variations
Synthetic students are stratified by the number of misconceptions held: 00, $1$, $2$, or $3+$.
D.3.7 Steering Experiments
Same dataset and model parameters as for Figure 6.
Training hyper-parameters.
[…]
  • $G=4$
  • Learning $\text{Learning rate}=1\times 10^{-4}$
  • $\beta=0.04$
  • Optimiser: Adam (adamw_torch, default settings)
  • LoRA configuration
    • $r=16$, $\alpha=16$
    • target_modules = ["q_proj", "k_proj", "v_proj", "o_proj", "gate_proj", "up_proj", "down_proj"]
    • Dropout: $0.05$
D.4 Prompts
Below are the different prompts used to query LLMs in our experiments.
1template: |
2 Please produce a concise summary of the following input with up to {max_words} words.
3 This summary should contain the information required to predict the students answer to any new question.
4 {cot_instruction}
5
6 Input:
7 {input_text}
8
9 Once again, please provide a **concise** summary of the students knowledge state.
10 {cot_instruction}
11 Keep your summary under {max_words} words.
Listing 1: Base Prompt for the Encoder.
1template: |
2 Here is some information about a student:
3 {bottleneck}
4
5 {y_prompt}
Listing 2: Base Prompt for the Decoder.
1template: |
2 Please produce a concise summary of the following input with up to {max_words} words.
3 This summary should contain the information required to predict the students answer to any new question.
4
5 Make sure to mention any misconception held by the student.
6 {cot_instruction}
7
8 Input:
9 {input_text}
10
11 Once again, please provide a **concise** summary of the students knowledge state.
12 {cot_instruction}
13 Keep your summary under {max_words} words.
Listing 3: Base Prompt for the Encoder when steering for mentioning "misconceptions" (Section B.4.1).
D.5 Code and hardware
Experiments were ran on four NVIDIA A100 GPUs with 80GB VRAM each, and 880GB of total RAM.

### v2

@@
-Language Bottleneck Models: A Framework for Interpretable Knowledge Tracing and Beyond
+Language Bottleneck Models for Qualitative Knowledge State Modeling
 Abstract
-Accurately assessing student knowledge is critical for effective education, yet traditional Knowledge Tracing (KT) methods rely on opaque latent embeddings, limiting interpretability.
-Even LLM-based approaches generate direct predictions or summaries that may hallucinate without any accuracy guarantees.
-We recast KT as an inverse problem: learning the minimum natural-language summary that makes past answers explainable and future answers predictable.
-Our Language Bottleneck Model (LBM) consists of an encoder LLM that writes an interpretable knowledge summary and a frozen decoder LLM that must reconstruct and predict student responses using only that summary text.
-By constraining all predictive information to pass through a short natural-language bottleneck, LBMs ensure that the summary contains accurate information while remaining human-interpretable.
-Experiments on synthetic arithmetic benchmarks and the large-scale Eedi dataset show that LBMs rival the accuracy of state-of-the-art KT and direct LLM methods while requiring orders-of-magnitude fewer student trajectories.
-We demonstrate that training the encoder with group-relative policy optimization, using downstream decoding accuracy as a reward signal, effectively improves summary quality.
+Accurately assessing student knowledge is central to education.
+Cognitive Diagnosis (CD) models estimate student proficiency at a fixed point in time, while Knowledge Tracing (KT) methods model evolving knowledge states to predict future performance.
+However, existing approaches either provide quantitative concept mastery estimates with limited expressivity (CD, probabilistic KT) or prioritize predictive accuracy at the cost of interpretability (deep learning KT).
+We propose Language Bottleneck Models (LBMs), where an encoder LLM produces textual knowledge state summaries, which a decoder LLM uses to predict future performance.
+This produces interpretable summaries that can express nuanced insights—such as misconceptions—that CD and KT models cannot capture.
+Extensive validation across synthetic and real-world datasets shows LBMs reveal qualitative insights beyond what CD and KT models can capture, while achieving competitive accuracy with improved sample efficiency.
+We demonstrate that the encoder and decoder can be fine-tuned with reinforcement learning and supervised fine-tuning respectively to improve both summary quality and predictive performance.
 1 Introduction
-The Challenge of Estimating a Student’s Knowledge State.
-A fundamental objective in education is accurately assessing what a student knows, identifying misconceptions, and understanding how their knowledge evolves over time posner1982accommodation ; larkin2012misconceptions ; chen2020impact .
+Knowledge state modeling
+A fundamental objective in education is accurately assessing what a student knows, identifying misconceptions, and understanding how their knowledge evolves over time (Posner et al., 1982; Larkin, 2012; Chen et al., 2020).
 Teachers intuitively achieve this through diagnostic reasoning: by observing students’ answers, they infer not merely correctness, but deeper patterns reflecting conceptual mastery or specific misunderstandings.
-For instance, a student consistently misapplying algebraic properties signals conceptual gaps requiring targeted intervention.
-Replicating this nuanced interpretability within AI would profoundly enhance personalized education at scale, augmenting human teaching by providing transparent, actionable insights into learners’ knowledge states (khosravi2022explainable, ).
-Building algorithms that can learn the knowledge state of students as they interact with educational material is an active field of study known as Knowledge Tracing (KT) (corbett_knowledge_1994, ; shen2024survey, ) (see Appendix A for a detailed introduction to KT).
-However, accurately modeling a student’s knowledge state in a way that is both predictive and interpretable remains a significant open challenge.
-Traditional KT methods (piech_deep_2015, ; zhang2017dynamic, ; pandey2019self, ; ghosh2020context, ) rely primarily on statistical, embedding-based representations.
-Such models lack transparency, representing knowledge states as high-dimensional latent vectors that provide limited interpretability or actionable insights for teachers and learners (Figure 1B).
-A truly impactful AI assistant in education must go beyond opaque predictions, explicitly articulating what the student understands and what they do not, in an accurate and interpretable manner.
-Limitations of Existing LLM-based approaches for KT.
-Recent approaches leveraging large language models (LLMs) have shown that LLMs can produce sensible predictions of students’ future behavior when provided with relevant information (li2024explainable, ; kim2024beyond, ), help mitigate the KT’s cold-start problem (lee2024language, ) and be finetuned to improve accuracy on KT tasks (wang2025llm, ).
-Nonetheless, these LLM-based methods still lack rigorous interpretability, as they either treat the model as a black-box or rely on free-form explanations that are susceptible to hallucination (bender2021dangers, ).
-Without ensuring faithful and grounded summaries, LLM-generated explanations fall short as reliable educational tools, emphasizing the need for a structured and principled foundation for generating interpretable, reliable, and adaptive student knowledge representations.
-Figure 1: Language Bottleneck Models for Knowledge Tracing.
-The observed questions and answer $\mathcal{D}=\{\mathcal{X},\mathcal{Y}\}$ are caused by a certain knowledge state $\mathcal{S}$ held by the student when answering questions.
-The language bottleneck aims to approximate this knowledge state in natural language ($\tilde{\mathcal{S}}$).
-Then, the predicted knowledge state is used to predict both the observed historical data $\mathcal{X}$ and future questions $\mathcal{Y}$.
-Knowledge Tracing as an Inverse Problem: Towards Concise and Interpretable Summaries.
-Existing Knowledge Tracing methods typically approach student modeling as a forward prediction task, using latent statistical embeddings to predict future responses.
-However, this approach provides limited insight into the underlying student knowledge.
-We reframe Knowledge Tracing as an inverse problem: viewing student responses as outcomes of an underlying knowledge state $\mathcal{S}$, manifested through the student’s cognitive process (Figure 1A).
-The goal is to recover an interpretable representation $\tilde{\mathcal{S}}$ of the student knowledge state.
-Crucially, answers to downstream questions are expected to arise from the same latent knowledge state—an assumption we discuss further in Section 2.2.
-A useful representation must therefore reconstruct observed answers, enable prediction of future responses, and remain interpretable and steerable for teachers.
-Introducing Language Bottleneck Models (LBMs).
-Motivated by the inverse-problem framing described above, we introduce Language Bottleneck Models (LBMs).
-As illustrated in Figure 1C, LBMs uses two Large Language Models (LLMs) as student knowledge state extractor (encoder) and interpreter (decoder), generating succinct textual summaries meant to capture the student knowledge state.
-Specifically, LBMs map sequences of student interactions $\mathcal{X}$ to a concise, natural-language summary $\tilde{\mathcal{S}}$.
-This summary then serves as the sole basis for reconstructing the observed responses $\mathcal{X}$ and accurately predicting future responses $\mathcal{Y}$.
-By design, this ensures that the generated explanations encompass precisely the information necessary to interpret past behavior and reliably forecast future student performance.
-Contributions and Paper Structure.
-We make three key contributions:
-  • We reframe knowledge tracing as an inverse problem and formalize a minimal-description objective that yields accurate and interpretable summaries of student knowledge states.
-  • We introduce Language Bottleneck Models (LBMs)—an encoder-decoder architecture that forces all predictive information through a concise natural-language summary.
-  • We empirically validate LBMs on synthetic and real-world datasets against strong KT and direct-prompting baselines, and show how reinforcement finetuning can train the encoder using downstream prediction accuracy.
-The remainder of this paper is organized as follows: Section 2 introduces the knowledge tracing problem and reframes it as an inverse problem.
-Section 3 presents our Language Bottleneck Model (LBM) framework, along with results motivating this approach, and training methods.
-Section 4 discusses related literature on knowledge tracing and interpretability.
-Section 5 presents experimental results validating the effectiveness of LBMs.
-Finally, Section 6 discusses the applicability of LBMs beyond knowledge tracing and concludes with directions for future research.
-2 Knowledge Tracing as an Inverse Problem
-In this section, we formally introduce the knowledge tracing problem.
-We then reframe it as an inverse problem—inferring interpretable knowledge states from observed responses, outlining key desiderata for such representations and the conditions under which this formulation is appropriate.
-2.1 Knowledge Tracing: Formal Problem Definition
-Knowledge Tracing (KT) is fundamentally a sequence modeling problem where the objective is to infer a student’s latent knowledge state based on observed interactions.
-Formally, let $\mathcal{X}=\{(q_{1},a_{1}),(q_{2},a_{2}),\ldots,(q_{t},a_{t})\}$ denote a sequence of question-answer pairs, where $q_{i}\in\mathcal{Q}$ represents a question from the question space $\mathcal{Q}$, and $a_{i}\in\mathcal{A}$ represents the student’s answer from the answer space $\mathcal{A}$.
-The goal of KT is to learn a function that predicts the probability of correctly answering a future question $q_{t+1}$ given the historical interaction sequence: $P(a_{t+1}|q_{t+1},\mathcal{X})$.
-While this formulation captures the predictive aspect of KT, it fails to address a crucial dimension: the interpretability.
-An ideal KT system should not only predict accurately but also provide educators with an explicit, human-understandable representation of the student’s knowledge state—such as mastered concepts, knowledge gaps and misconceptions.
-The importance of interpretability for KT is further discussed in the Appendix.
-2.2 Knowledge Tracing as an Inverse Problem: From Black–Box Predictions to Interpretable Summaries
-The conventional KT formulation (§2.1) focuses on predictive accuracy but offers limited insight into the underlying knowledge that drives student behavior.
-We instead frame KT as an inverse problem: in many practical scenarios—such as a quiz completed in a single sitting—a student’s responses can largely be explained by a fixed underlying knowledge state, expressed through the student’s cognitive process of answering questions.
-Most observed behavior reflects this latent state, with occasional deviations due to noise—such as slips (careless mistakes) or guesses (lucky correct answers), as formalized in Bayesian KT corbett_knowledge_1994 .
-The goal, then, is to recover an interpretable representation $\tilde{S}$ of the knowledge state that accounts for the observed interactions and supports accurate future predictions.
-Framing KT in this way highlights three desiderata.
-(1) Self-consistency: the inferred summary $\tilde{S}$ should reconstruct the past by making the observed answers in $\mathbf{X}$ likely under a decoder.
-(2) Predictive sufficiency: $\tilde{S}$ must retain all information necessary to accurately predict future responses to unseen questions.
-(3) Human interpretability and steerability: the summary should be readable and editable by domain experts, enabling the diagnosis of student misconceptions and manual refinement or augmentation of $\tilde{S}$ to guide future predictions.
-When is the inverse formulation appropriate?
-Although real learning is dynamic, many high–stakes diagnostic situations (end–of–unit quizzes, placement tests, or short tutoring sessions) span minutes rather than weeks.
-In such windows, it is reasonable to assume a quasi–static knowledge state.
-In this work, we explicitly focus on scenarios characterized by such quasi–static knowledge states.
-For longer horizons, the knowledge state might be better modeled as piecewise-static or smoothly evolving $S(t)$.
-We leave the extension of our framework to dynamically evolving knowledge states for future work.
+Limitations of Cognitive Diagnosis and Knowledge Tracing
+Cognitive Diagnosis (CD) (Templin et al., 2010; Wang et al., 2024) models produce diagnostic reports but are constrained to proficiency estimates over a fixed set of knowledge concepts.
+In parallel, Knowledge Tracing (KT) (Shen et al., 2024) models excel at predicting future performance based on observed past responses, yet remain limited to either estimating mastery over knowledge concepts (Corbett and Anderson, 1994; Zhou et al., 2024) or operating as black-box models requiring post-hoc interpretability (e.g.
+DKT (Piech et al., 2015; Ghosh et al., 2020)).
+Limitations of existing LLM-based approaches
+Recent approaches leveraging large language models (LLMs) have shown that LLMs can produce sensible predictions of students’ future behavior when provided with relevant information (Li et al., 2024a; Kim et al., 2024), help mitigate the KT’s cold-start problem (Lee et al., 2024) and be finetuned to improve accuracy on KT tasks (Wang et al., 2025).
+Nonetheless, these LLM-based methods still lack rigorous interpretability, as they either treat the model as a black-box or rely on free-form explanations that are susceptible to hallucination (Bender et al., 2021).
+As a result, such methods fail to provide grounded, reliable representations of knowledge that can be trusted in educational practice.
+This gap motivates a principled approach where interpretability is an intrinsic design constraint.
+Figure 1: Language Bottleneck Models for Knowledge Modeling.
+(A) Past and future behavior $\mathcal{X}$ and $\mathcal{Y}$ are caused by a certain knowledge state $\mathcal{S}$ held by the student when answering questions.
+(B) CD and KT models represent the knowledge state via quantitative proficiency vectors or opaque latent embeddings.
+(C) LBMs approximate the knowledge state using natural language summaries.
+The example knowledge states shown are taken from the case-study presented in Section 4.1.
+Knowledge state modeling as an inverse problem
+Assessing student knowledge can be framed as an inverse problem (Figure 1A): observable answers are generated by an underlying knowledge state $\mathcal{S}$, and the goal is to infer a faithful approximation $\tilde{\mathcal{S}}$.
+This formulation is commonly done in Cognitive Diagnosis, which constrains $\tilde{\mathcal{S}}$ to fixed quantitative proficiency estimates over knowledge concepts (Figure 1B), providing interpretability but lacking flexibility to capture nuanced knowledge states.
+In this work, we adopt an inverse-problem framing in which $\tilde{\mathcal{S}}$ takes the form of concise, free-form natural language.
+This shifts the knowledge state from proficiency estimates over knowledge concepts to a flexible description that can capture nuanced insights such as misconceptions.
+Language Bottleneck Models
+We introduce Language Bottleneck Models (LBMs), a modeling framework that uses natural language as an interpretable intermediate representation.
+Constraining all predictive information to flow through a textual bottleneck ensures that the model’s predictions are grounded in human-readable descriptions.
+In our knowledge state modeling setting (Figure 1C), an encoder LLM summarizes a student’s interaction history $\mathcal{X}$ into natural language $\tilde{\mathcal{S}}$, while a decoder LLM reconstructs past responses and predicts future ones using only that summary.
+This constitutes a language-based approach to Cognitive Diagnosis that captures nuanced insights beyond quantitative proficiency scores while maintaining strong predictive performance.
+Contributions.
+  • We cast knowledge state modeling as an inverse problem over open-ended textual representations—building on ideas from Cognitive Diagnosis but replacing quantitative concept proficiencies with a flexible natural-language knowledge state.
+  • To instantiate this, we introduce Language Bottleneck Models (LBMs), which encode observed student behavior into predictive text-based summaries of knowledge states.
+  • We present a rigorous qualitative evaluation using LLM-as-a-judge metrics and a detailed case study, demonstrating that LBMs provide nuanced insights—such as misconceptions—beyond what CD and KT models can capture.
+  • We evaluate LBMs against 14 KT and CD baselines on three datasets, showing competitive accuracy with superior sample efficiency, and demonstrate that fine-tuning the encoder and decoder enhances predictive performance.
+2 Knowledge State Modeling as an Inverse Problem over Textual Representations
+2.1 Preliminaries and Notation
+We consider data consisting of student interactions with educational questions.
+At each time step $t$, a student is presented with a question $q_{t}\in\mathcal{Q}$ and optionally knowledge concept (KC) information $k_{t}\in\mathcal{K}$ and provides a response $r_{t}\in\mathcal{R}$, which is evaluated for correctness $c_{t}\in\{0,1\}$.
+We represent an interaction as $x_{t}=(q_{t},k_{t},r_{t},c_{t})$, and a student’s interaction history up to time $t$ as $H_{t}=(x_{1},\ldots,x_{t})$.
+The standard predictive task, often referred to as Knowledge Tracing (KT) (Corbett and Anderson, 1994; Shen et al., 2024), is to estimate $p(c_{t+1}\mid q_{t+1},H_{t})$, the probability that the student will answer a new question $q_{t+1}$ correctly, conditioned on their interaction history.
+This task definition captures the forward-prediction aspect of knowledge state modeling, but does not yet address how the underlying knowledge that drives these responses should be represented.
+2.2 Knowledge State Modeling as an Inverse Problem
+An alternative view is to frame knowledge state modeling as an inverse problem: observed responses are generated by a latent knowledge state $\mathcal{S}$ through the student’s cognitive process, and the goal is to recover an approximation $\tilde{\mathcal{S}}$ of this state from the observed responses.
+This perspective is central to Cognitive Diagnosis (CD) models (Reckase, 2006; De La Torre, 2009; Templin et al., 2010; Wang et al., 2022, 2024), which produce diagnostic reports of concept mastery from observed responses.
+However, CD typically restricts $\tilde{\mathcal{S}}$ to quantitative mastery or proficiency vectors based on predefined or inferred concepts, limiting expressivity.
+Meanwhile, deep learning-based KT methods prioritize predictive accuracy without explicitly recovering $\tilde{\mathcal{S}}$, representing knowledge states as high-dimensional embeddings that lack transparency (Piech et al., 2015; Zhang et al., 2017; Pandey and Karypis, 2019; Ghosh et al., 2020).
+Bayesian KT (Corbett and Anderson, 1994) as well as recent KT extensions introduce diagnostic reports similar in spirit to CD, or use interpretable latent states (Yeung, 2019; Minn et al., 2022; Chen et al., 2023; Park et al., 2024; Zhou et al., 2024).
+However these approaches remain bound to estimates of KC mastery or rely on post-hoc interpretability.
+Key assumption: constant knowledge state A practical assumption underlying this formulation is that a knowledge state can be treated as constant within short diagnostic windows (e.g., unit tests, placement exams, or tutoring sessions).
+This aligns with CD models (see §2.1 in Wang et al.
+(2024)), and distinguishes them from KT approaches which model the evolution of $\mathcal{S}_{t}$ across longer periods where the underlying knowledge state is expected to change.
 2.3 Natural Language as the Interface
-While a student’s knowledge state is inherently latent, natural language offers a flexible and interpretable medium for expressing it.
-Unlike fixed concept vocabularies, natural language accommodates open-ended reasoning patterns, misconceptions, and levels of mastery.
-In this work, we propose to use concise natural language summaries as proxies for student knowledge, designed to capture the latent cognitive state in a way that is both human-readable and predictive.
-The next section introduces a model family that operationalizes this idea.
+Our formulation follows CD approaches in adopting the inverse problem framing of inferring a diagnostic report from observed responses, but instead of restricting $\tilde{\mathcal{S}}$ to quantitative mastery scores, we model it through concise textual summaries.
+Natural language provides an interpretable and expressive medium—capable of describing arbitrary reasoning patterns or misconceptions, a key focus in education research (Smith III et al., 1994; Wang et al., 2020; King et al., 2024)—while remaining human-understandable.
+In the next section, we introduce Language Bottleneck Models (LBMs), which operationalizes this idea by compressing student interaction histories into concise textual representations that preserve predictive information.
 3 Language Bottleneck Models
-Figure 2: Zero-shot knowledge state encoding compared to ground truth.
-LLM models with different capabilities are prompted to write a summary of the knowledge state (top left panel) of a student, given 100 observed questions and answers provided in text.
+3.1 Formal Definition
+We propose Language Bottleneck Models (LBMs) for Knowledge State Modeling via textual summaries: an LLM-based, two-stage architecture designed to infer a predictive text-based knowledge state from a student’s interaction history.
+[…]
+In later sections, we describe how both the encoder $f_{\theta}$ and decoder $g_{\phi}$ can be trained.
+However, we first show with the following motivating experiments that encoding is a harder task than decoding: when given high-quality knowledge state summaries, strong off-the-shelf LLMs can achieve near-perfect prediction accuracy.
+3.2 Motivating Observations
+We motivate the design of LBMs by two observations.
+Observation 1: Given a good knowledge state summary, strong LLMs can decode with high fidelity.
+To test this in an idealized setting, we used a Synthetic dataset where each student’s knowledge state and their resulting answers to questions are programmatically generated.
+Figure 2 evaluates the performance of different decoder models when given direct access to this perfect, "ground-truth" summary of the student’s latent knowledge (example knowledge state and summaries are shown in Figure A1 in the Appendix, and full dataset details are in §4).
+Stronger models like GPT-4o achieve nearly perfect accuracy ($98\%$), indicating that the bottleneck representation is indeed sufficient to drive effective downstream prediction for closed-form questions—provided it captures the right information.
+Figure 2: Accuracy ± SEM (N=200) on Synthetic dataset given ground truth knowledge state summaries.
+Observation 2: Summarizing knowledge states from raw interactions is non-trivial.
+Standard LLM summarization approaches can capture high-level skill mastery but often fail to identify crucial latent patterns like student misconceptions (see an example comparison on our Synthetic dataset Figure A1 in the Appendix).
+Together, these observations suggest that the key challenge lies in learning an encoder that produces faithful summaries, rather than in the decoding step itself.
+3.3 Training Objective and Optimization
+3.3.1 Encoder
+We propose a reinforcement learning-based approach to train an encoder to produce more faithful and predictive summaries by using downstream decoder accuracy as reward.
+Summary generation and prediction Given an interaction history $H_{t}=(x_{1},\ldots,x_{t})$, the encoder $f_{\theta}$ maps a subset of interactions $\mathcal{X}_{\text{enc}}\subseteq H_{t}$ to a textual summary $\tilde{\mathcal{S}}=f_{\theta}(\mathcal{X}_{\text{enc}})$.
+The decoder $g$ then conditions on $\tilde{\mathcal{S}}$ to predict responses for two sets $\mathcal{X}$ and $\mathcal{Y}$: interactions used for reconstruction and prediction, respectively.
+In practice, $\mathcal{X}$ and $\mathcal{Y}$ may be chosen flexibly to include held-out past responses, future responses, or both.
+Reward function Given predicted interactions $\tilde{\mathcal{X}}=\{g(q,\tilde{\mathcal{S}}),q\in\mathcal{X}\}$ and $\tilde{\mathcal{Y}}=\{g(q,\tilde{\mathcal{S}}),q\in\mathcal{Y}\}$, the reward for a summary $\tilde{\mathcal{S}}$ is defined as
+$),R(\tilde{\mathcal{S}};g)=\phi\!\Big(\texttt{Acc}\big(\tilde{\mathcal{X}},\mathcal{X}\big),\texttt{Acc}\big(\tilde{\mathcal{Y}},\mathcal{Y}\big),\;|\tilde{\mathcal{S}}|,\;\Omega(\tilde{\mathcal{S}}$$\Big)$, (1)
+[…]
+Optimization via GRPO We optimize the encoder $f_{\theta}$ with Group Relative Policy Optimization (GRPO) (Shao et al., 2024).
+For each input $\mathcal{X}_{\text{enc}}$, the encoder generates $G$ candidate summaries $\{\tilde{\mathcal{S}}^{1},\ldots,\tilde{\mathcal{S}}^{G}\}$, each receiving reward $r_{i}=R(\tilde{\mathcal{S}}^{i};g)$.
+We then compute group-relative advantage $\hat{A}_{i}=\frac{r_{i}-\text{avg}(\textbf{r})}{\text{std}(\textbf{r})}$ and train with the DAPO loss (Yu et al., 2025):
+$\mathcal{L}(\theta)=-\frac{1}{\sum_{i=1}^{G}|\tilde{\mathcal{S}}^{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|\tilde{\mathcal{S}}^{i}|}\biggr[\frac{\pi_{\theta}(\tilde{\mathcal{S}}^{i}_{t}|\mathcal{X}_{\text{enc}},\tilde{\mathcal{S}}^{i}_{<t})}{[\pi_{\theta}(\tilde{\mathcal{S}}^{i}_{t}|\mathcal{X}_{\text{enc}},\tilde{\mathcal{S}}^{i}_{<t})]_{\text{no grad}}}\hat{A}_{i}\\ -\beta\mathbb{D}_{\mathrm{KL}}[\pi_{\theta}||\pi_{\text{ref}}]\biggr]$ (2)
+where $\beta$ is a hyperparameter and $\pi_{\text{ref}}$ is the original model.
+3.3.2 Decoder
+The decoder $g_{\phi}$ is trained via standard supervised fine-tuning (SFT).
+We generate knowledge state summaries $\tilde{\mathcal{S}}$ for a set of training students and pair them with ground-truth correctness labels for target questions $q\in\mathcal{Y}$ .
+The model is optimized to predict the correct binary outcome (e.g., "Yes" or "No") conditioned on the summary $\tilde{\mathcal{S}}$ and query question.
+3.4 Steerability of the Estimated Knowledge State
+The natural-language summaries generated by LBMs allow for various human-model interactions (detailed in Appendix D).
+(1) Prompt engineering the encoder.
+Since the encoder $f_{\theta}$ is itself an LLM, its behavior can be shaped through prompt design, such as system instructions or in-context examples (Brown et al., 2020).
+(2) Steering via reward signals.
+Rewards to steer the encoder towards human preferences can be incorporated through the $\Omega(S)$ term in Eq.
+1.
+(3) Augmenting with student-specific information.
+Educators can supplement the model with additional knowledge not present in observed data—either by augmenting encoder inputs or by editing the generated summary before decoding.
+This enables integration of recent classroom observations or specific misconceptions identified through in-person interactions.
+4 Experiments
+Datasets
+We evaluate LBMs and baseline models on a synthetic arithmetic benchmark and two real-world datasets (Table 4).
+Our Synthetic dataset (Appendix  B.4.1) simulates learners answering addition, subtraction, multiplication, and division questions; each student is assigned mastered skills, unmastered skills, and systematic misconceptions.
+We filter Eedi and XES3G5M for single-session trajectories (max inter-question gaps of 3 and 10 minutes respectively) with at least 40 and 34 interactions respectively to approximate the static knowledge state assumption common in CD (Wang et al., 2024).
+We evaluate XES3G5M in both Chinese and English using translations from Ozyurt et al.
+(2024) (see Appendix B.4).
+Table 1: Overview of datasets.
+AVG#log and STD#log>1 are defined following  Wang et al.
+(2022) as respectively the average number of logs per student per KC, and the mean standard deviation of score per student and per KC.
+Dataset Synthetic Eedi (Filt.) XES3G5M (Filt.)
+#Students 2,000 623 996
+#Questions 5,000 3,401 3,221
+#KCs 4 1,284 803
+#Logs 420,000 28,947 57,788
+#Logs/stud. 210 $\geq$40 $\geq$34
+[…]
+[…]
+[…]
+Figure 3: Case study: comparing CD and LBM knowledge states.
+Given a student from the Synthetic dataset, we compare proficiency estimates across knowledge concepts (KCs) obtained from a trained NeuralCDM model to the text-based knowledge state generated by a trained LBM model.
+Models We evaluate LBMs across LLM backbones of different sizes and capabilities, both open-source (Qwen 2.5 3B and 7B (Team, 2024), Gemma 3 12B and 27B (Team, 2025)), and closed-source (GPT-4o-mini, GPT-4o and GPT-5 (Achiam et al., 2023)).
+We train one decoder model by finetuning a Gemma3-12B model with LoRA using on a GPT-5 encoder (labeled gpt-5+gem-12b* in the results, see details in § B.2).
+Unless noted otherwise, the same backbone LLM is used for both the encoder and decoder, and prompt all models to provide their response directly without chain-of-thought.
+We run GPT-5 with reasoning_effort=minimal configuration.
+Hyper-parameters and prompt templates are provided in Appendix B.
+Baselines We compare LBMs against 9 Knowledge Tracing methods: DKT (Piech et al., 2015), DKVMN (Zhang et al., 2017), SAKT (Pandey and Karypis, 2019), AKT (Ghosh et al., 2020), Deep IRT (Yeung, 2019), SAINT (Choi et al., 2020), SimpleKT (Liu et al., 2023a), QIKT (Chen et al., 2023), GKT (Nakagawa et al., 2019)) implemented with the pyKT library (Liu et al., 2022) and 5 Cognitive Diagnosis methods (IRT (Rasch, 1993), MIRT (Reckase, 2006), DINA (Junker and Sijtsma, 2001), KaNCD and NeuralCDM (Wang et al., 2022)) implemented with the EduCDM library (bigdata-ustc, 2021).
+We also run each LLM via direct prompting, where the LLM predicts answers from the full interaction history without a bottleneck.
+Training details for all baselines are given in Appendix B.
+Systematic evaluation of knowledge state summaries We systematically evaluate the quality of the knowledge state summaries produced by different encoder models on the Synthetic dataset using LLM-as-a-judge (Li et al., 2025).
+Specifically we prompt a GPT-5 model to compare each summary to the corresponding ground-truth knowledge state across the following dimensions (see Appendix C for full details and prompts):
+  • Global score: overall alignment with the true state,
+  • Construct accuracy: correctly assessed constructs,
+  • Misconception detection: correctly identified student misconceptions,
+  • Misconception false positives: misconceptions mentioned but absent in the ground truth,
+  • Specificity: precision vs. vagueness of the description,
+  • Confidence calibration: degree of over- or under-confidence.
+Figure 4: Systematic evaluation of the summaries produced by different encoder LLMs over 200 test students from the Synthetic dataset: (Left) Average construct mastery accuracy across constructs (%) and average overall score (1-5 scale); (Middle) Average misconception detection rate (%) and average number of misconception false positives per summary (count); (Right) Confidence Index (1-3, 1=under-confident, 2=appropriately calibrated, 3=over-confident) and Specificity Index (1-3, higher is more specific).
+Error bars represent the standard error (N=200).
+4.1 Qualitative Insights
+4.1.1 Case-study: comparing CD and LBM knowledge state representations
+A key advantage of LBMs over CD models is the ability to capture nuanced insights about the student knowledge state, such as misconceptions.
+We illustrate this with a case-study Figure 3.
+We train a state-of-the-art CD model (NeuralCDM) trained on our Synthetic dataset, which achieves strong predictive performance (AUC: 0.96, Acc: 0.90 on the test set).
+From this model, we extract proficiency estimates across knowledge concepts using the learned student embedding vector, producing a typical CD diagnostic report.
+We then compare this to the knowledge state summary generated for the same student by an LBM (trained Gemma-12B encoder with frozen Gemma-27B decoder, as in §4.2).
+While NeuralCDM reliably captures general KC proficiency, its estimates are influenced by misconceptions without explicitly identifying them.
+In contrast, LBMs capture overall proficiency and uncover specific misconceptions (e.g., errors with negative numbers or with operand-6).
+This ability to provide nuanced, qualitative insights into student knowledge states sets LBMs apart from CD methods.
+4.1.2 Systematic evaluation of knowledge summaries
+To further assess the interpretability of the knowledge states produced by LBMs, we systematically compare their summaries to ground truth knowledge states from the Synthetic dataset using GPT-5 as LLM-as-a-judge (Li et al., 2025).
+The results are shown in Figure 4.
+Figure 4a shows the overall alignment with the ground truth summary (score from 1=’mostly incorrect’ to 5=’strongly aligned’), and the construct mastery accuracy.
+Model capability correlates with performance on both metrics, with GPT-5 achieving the highest scores.
+Figure 4b plots the misconception-detection rate against the misconception false positive rate.
+Qwen2.5 and GPT-4o/4o-mini exhibit lower hallucination but also lower detection; Gemma-3 detects more misconceptions but at the cost of more false positives; GPT-5 outperforms all models on both dimensions.
+Figure 4c evaluates summary specificity and confidence calibration.
+Gemma-3 models produce more specific summaries than Qwen2.5, but both families tend to be over-confident relative to GPT models, with GPT-5 displaying the best overall specificity and calibration.
+Refer to captionA
+Refer to captionB
+Figure 5: Training the encoder on Synthetic dataset.
+(A) Evolution of the test accuracy during training of a Gemma3-12B encoder, with a frozen Gemma3-27B decoder.
+Trained on $800$ students and tested on $200$ students, with $|\mathcal{X}|=50$ questions per input trajectory and $|\mathcal{Y}|=20$ questions to predict per student.
+Bottom row shows the evolution of generated summaries over the course of training for an example student.
+Text is colored green (exact), orange (approximate), or red (false) based on ground-truth.
+(B) Systematic evaluation of generated summaries over encoder training.
+Evaluation axes are the same as Figure 4(a) and (c).
+Confidence intervals show the standard error (N=200).
+4.2 Training the Encoder
+As outlined in §3.3.1, downstream accuracy can serve as a reward signal to train an encoder to produce increasingly accurate summaries.
+We demonstrate this by training a Gemma3-12B encoder using a frozen Gemma3-27B decoder on 800 students.
+We set the reward as the decoder accuracy across $|\mathcal{Y}|=20$ unseen questions $R(\tilde{\mathcal{S}};g)=\texttt{Acc}\big(\tilde{\mathcal{Y}},\mathcal{Y}\big)$, train with a LoRA adapter (Hu et al., 2022) and evaluate on 200 unseen test students.
+Effect on accuracy Figure 5 shows the encoder progressively improving summary quality and quickly outperforming GPT-4o.
+The figure illustrates this through an example student who mastered all constructs except subtraction and fails any multiplication involving 9.
+The initial summary contains inaccuracies and misses this systematic misconception, while the final summary successfully captures the student’s complete knowledge state.
+Stratifying by knowledge state complexity, we observe larger gains for more complex cases (Appendix A.3.1).
+Qualitative evolution of summaries We use the same LLM-as-a-judge approach as in the previous section to systematically evaluate the generated summaries over training.
+Figure 5 shows that the summaries gradually better capture concept mastery and align more with the ground truth knowledge state over training.
+Moreover, as summaries also become more specific and better calibrated.
+Misconception detection and false positive rates do not significantly change over the course of training (Appendix A.3.2).
+4.3 LBM vs Knowledge Tracing Methods
+We present comprehensive results comparing LBMs to baseline methods across all datasets (detailed results in Table A.2 in the Appendix).
+Here we highlight key findings and insights from these experiments.
+Refer to captionA
+Refer to captionB
+Figure 6: (Left) Accuracy of LBM vs the top-2 KT and CD models across datasets.
+Models are grouped by LBMs (blue), KT models (orange) and CD models (red).
+Top 2 KT and CD models are selected based on average accuracy across all three datasets.
+LBMs are evaluated zero-shot except gpt-5+gem-12b* where the Gemma3-12b decoder is trained.
+Full results for other models are available Table A.2 in the Appendix.
+(Right) Accuracy of the best LBM, KT and CD models on the Synthetic dataset.
+The x axis shows the total number of question seen by the model, i.e.
+"Traj.
+Length" x "# Students".
+Note that $\text{Students}=1$ for LBMs as they are evaluated zero-shot.
+Results are averaged across N=10 runs for KT and CD models and N=3 for LBMs, with error bars showing the standard error.
+Performance Figure 6 compares zero-shot and trained LBMs against the top KT and CD models (full results in Table A.2).
+Zero-shot LBMs with powerful backbones (GPT-4o, GPT-5) rival the performance of the best baselines across all datasets.
+Training the decoder further improves accuracy: a fine-tuned Gemma-12B decoder outperforms the zero-shot GPT-5 decoder when both use the same GPT-5 encoder, and further closes the gap in accuracy with KT/CD models.
+Notably, this "GPT-5 Encoder+trained Gemma3-12b Decoder" configuration ranks first in AUC among all KT/CD models on Eedi and third on XES3G5M-English.
+(Table A.2).
+Sample efficiency Figure 6B compares the performance of the best zero-shot LBM model (GPT-5) to the best CD and KT models (KaNCD and AKT) on the Synthetic dataset.
+LBMs with a GPT-5 backbone achieves comparable accuracy to KT/CD methods with significantly less training data.
+Since CD/KT methods rely on statistical patterns, they require substantially more observations before reaching strong predictive power, while LBMs demonstrate strong zero-shot performance.
+4.4 Steering LBM behavior
+We demonstrate multiple steering strategies described in §3.4.
+Providing explicit misconception information during encoder training produces substantially stronger learning effects than adding the same information at the decoder stage (Appendix A.4.1), suggesting that the encoder uses this additional context to better interpret patterns in student responses.
+We also show the encoder can be steered during training to explicitly mention misconceptions through reward signals (Appendix A.4.2), and illustrate how supplementing the summary with information not present in the input can significantly improve decoder accuracy (Appendix A.4.3).
+4.5 Ablation Experiments
+How much information is lost by the bottleneck?
+Table 2 compares the accuracy of LBM models to directly predicting new questions from the observed student data.
+Despite the information bottleneck, LBM accuracy typically remains within $2\%$ of direct prediction—and often surpasses it.
+Figure 7 shows that this gap decreases for longer bottleneck token limits, highlighting a trade-off between conciseness and predictive accuracy.
+Which of the encoder or decoder is most critical for LBMs?
+We evaluate LBMs with different encoder–decoder pairings (Appendix A.5.1).
+Using a strong model (GPT-4o) as the encoder with weaker models as decoders yields accuracies $5-10\%$ higher than when the stronger model is used as the decoder.
+This confirms our hypothesis that extracting relevant information for the summary (encoding) is more challenging than predicting future answers given a summary (decoding).
+Do LBMs require knowledge concept information?
+Table LABEL:app:tab:ablation compares LBMs with and without KC information in the input prompt on the Synthetic and EEDI (Filtered) datasets.
+Performance does not significantly change, demonstrating that LBMs do not fundamentally require KC information.
+Figure 7: Evolution of LBM accuracy with bottleneck length (GPT-4o backbone).
+Table 2: Accuracy results for Direct and LBM methods. Bold indicates models for which the LBM accuracy is no more than $2\%$ below the Direct baseline (Welch’s t-test, details in Appendix B.5.1).
+synthetic eedi Direct LBM $\Delta$ Direct LBM $\Delta$ Qwen2.5-3B .56 ± .00 .61 ± .01 +.06 .35 ± .00 .38 ± .00 +.03 Qwen2.5-7B .64 ± .00 .65 ± .01 +.01 .65 ± .00 .58 ± .01 -.07 gemma-3-12b .63 ± .00 .79 ± .00 +.17 .58 ± .00 .62 ± .01 +.04 gemma-3-27b .62 ± .00 .78 ± .01 +.16 .67 ± .00 .65 ± .01 -.02 gpt-4o-mini .81 ± .01 .78 ± .01 -.03 .67 ± .01 .61 ± .02 -.05 gpt-4o .85 ± .01 .80 ± .01 -.05 .66 ± .00 .66 ± .02 +.00 gpt-5 .87 ± .00 .87 ± .01 +.00 .71 ± .02 .68 ± .01 -.03
+5 Related Work
+We review related works from the Cognitive Diagnosis and Knowledge Tracing literature, as well as concept bottleneck models.
+See Appendix F for an extended review of related works, and Table F1 for a high-level comparison of LBMs with CD and KT.
+Cognitive Diagnosis Cognitive Diagnosis Models (CDMs) infer student knowledge states from observed responses.
+Classical approaches include Item Response Theory (IRT) and Multidimensional IRT which measure continuous proficiency scores (Rasch, 1993; Reckase, 2006), and the DINA model and its variants which estimate binary mastery of knowledge concepts (De La Torre, 2009).
+Recent deep learning variants like NeuralCDM (Wang et al., 2022) and RCD (Gao et al., 2021) use neural networks and graph architectures to model complex relationships between students, questions, and knowledge concepts.
+However, these models provide only quantitative skill mastery estimates over knowledge concepts.
+Knowledge Tracing Knowledge Tracing methods model student learning to predict future performance.
+Deep learning approaches like DKT (Piech et al., 2015), DKVMN (Zhang et al., 2017) and AKT (Ghosh et al., 2020) employ neural architectures.
+Despite strong predictive performance, these models represent knowledge as abstract latent vectors, limiting interpretability.
+Several recent works have proposed more interpretable KT.
+Early Bayesian approaches (Corbett and Anderson, 1994; Käser et al., 2017) and recent probabilistic models (Minn et al., 2022; Zhou et al., 2024) learn interpretable latent states representing student proficiency over knowledge concepts, while IRT-based methods (Yeung, 2019; Chen et al., 2023) combine deep learning with item response theory for meaningful latent representations.
+Other works improve interpretability through learned question relationships (Tong et al., 2022), explainable subsequences (Li et al., 2023), or option tracing (Ghosh et al., 2021).
+However, these approaches remain fundamentally constrained to quantitative concept proficiency estimation or require post-hoc interpretability.
+Finally, recent LLM-based approaches have been proposed for knowledge tracing tasks (Li et al., 2024a; Wang et al., 2025), but they generally remain opaque, directly predicting future answers without enforcing interpretable intermediate representations.
+Concept Bottleneck Models Concept Bottleneck Models (CBMs) (Koh et al., 2020) improve interpretability by using human-understandable concept activations as intermediates between inputs and predictions.
+However, CBMs typically rely on finite predefined concept sets, limiting applicability to complex tasks like knowledge tracing.
+Recently, Explanation Bottleneck Models (XBMs) (Yamaguchi and Nishida, 2024) use textual rationales as intermediates for vision classification.
+While our Language Bottleneck Models (LBMs) adopt this language bottleneck concept, they differ fundamentally: unlike XBMs’ instance-specific rationales, LBM summaries capture implicit knowledge states that generalize to future, unknown questions, requiring holistic, adaptable summaries rather than task-specific rationales.
+6 Discussion
+Why can’t we just prompt GPT-4o directly?
+The split encoder-decoder architecture of LBMs offers three key advantages over direct LLM prompting: it creates a global student model with a single latent summary shared across all predictions rather than isolated per-question reasoning; it ensures faithful summaries through a closed-loop decoding objective that penalizes non-predictive summaries; and it provides an explicit interface layer that teachers can read, steer and intervene on.
+Wider applicability LBMs extend beyond education to any task requiring compact, human-readable summaries with predictive power.
+The minimal ingredients needed are: (1) a sequence of observations about an entity, (2) a need to predict future behaviors of that entity, and (3) value in having interpretable representations.
+For example, clinical decision support could distill patient data into textual state descriptions that forecast outcomes while remaining auditable; preventive maintenance could compress sensor logs into explanations predicting machine failure; customer success teams could summarize interaction histories to forecast churn.
+Limitations and Future Work LBMs face several constraints including context length limitations, requirements for textual question content, and substantial computational costs.
+Future extensions could address these through iterative encoding for longer inputs, active sensing for optimal question selection, adaptation for evolving knowledge states, expansion beyond question-answer data, and integration with pedagogical techniques like LearnLM (Team et al., 2024).
+These limitations and extensions are discussed in detail in Appendix G.
+Appendix
+Appendix A Extended Results
+A.1 Motivating Observation 2
+Figure A1 compares summaries produced by different LLMs when prompted to describe a student’s knowledge state from 50 question-answer pairs from a synthetic dataset (see Section 4 for details).
+While all models capture high-level skill mastery, only one correctly identifies a misconception (errors with negative numbers) out of the four existing ones, illustrating that standard summarization approaches often miss crucial latent patterns.
+Figure A1: Zero-shot knowledge state encoding compared to ground truth.
+LLM models with different capabilities are prompted to write a summary of the knowledge state (top left panel) of a student, given 50 observed questions and answers provided in text.
 The ground truth knowledge state (top right panel) describing the student behavior has three main components: constructs mastered, constructs not mastered, and misconceptions.
 All three models capture the construct mastery correctly, but are not able to capture any misconception, beside o1 which notices the negative numbers misconception (bottom row).
-Bottom right symbols correspond to components in Figure 1.
-3.1 Formal Definition
-To tackle the inverse problem formulation of KT introduced in Section 2, we propose Language Bottleneck Models (LBMs): a two-stage architecture for inferring interpretable and accurate knowledge states from a student’s past interactions.
-Let $X_{s}\subseteq X$ denote a subset of observed student responses, and let $Y_{s}\subseteq Y$ denote a set of future questions whose answers the model aims to predict. An LBM is defined by:
-  • An encoder $f_{\theta}:X_{s}\rightarrow S$, which maps a history of interactions to a textual summary $S$ of the student’s implicit knowledge state.
-  • A decoder $g:S\rightarrow\hat{Y}_{s}$, which uses only the textual summary to predict the student’s answers to new questions.
-The explicit use of natural language as a bottleneck encourages the model to produce summaries that balance brevity with completeness.
-If the language summary omits essential details, predictions of future responses will suffer; conversely, if the summary is overly detailed or redundant, its utility for educators decreases.
-Thus, this design ensures the resulting textual summary is simultaneously concise, informative, and directly interpretable by humans.
-The following sections motivate this approach and outline how LBMs can be trained effectively.
-3.2 Motivating Observations
-Before detailing our training procedure, we highlight two motivating observations that justify the LBM framework and its constraints.
-Observation 1: Summarizing a knowledge state is challenging.
-Figure 2 compares summaries produced by different LLMs when prompted to describe a student’s knowledge state from 100 question-answer pairs from a synthetic dataset (see Section 5 for details).
-While all models capture high-level skill mastery, only one correctly identifies a misconception (errors with negative numbers) out of the four existing ones, illustrating that standard summarization approaches often miss crucial latent patterns.
-Figure 3: Accuracy on synthetic dataset given ground truth knowledge state summaries.
-Observation 2: Given a good summary, capable LLMs can decode with high fidelity.
-Figure 3 evaluates the performance of different decoder models when given access to the ground truth summary of the student’s latent knowledge state (an example of a ground-truth summary is shown in Figure 2).
-Stronger models like GPT-4o achieve nearly perfect accuracy ($98\%$), indicating that the bottleneck representation is indeed sufficient to drive effective downstream prediction—provided it encodes the right information.
-Together, these observations motivate the inverse formulation: while generating summaries from raw interactions is hard, once obtained, these summaries are powerful inputs for reasoning and prediction.
-This validates our approach of learning encoders that optimize predictive utility via interpretable, compressed natural language representations.
-3.3 Training Objective and Optimization
-We train LBMs to produce informative and interpretable summaries of student knowledge states.
-As demonstrated above, strong decoders such as GPT-4o can achieve near-perfect accuracy when provided with accurate knowledge summaries.
-We therefore optimize the encoder $f_{\theta}$ to produce such summaries, while freezing the decoder $g$ to a strong LLM during training.
-Summary generation and evaluation.
-Given a student’s interaction history $X$, we sample a subset $X_{\text{enc}}\subseteq X$ of $N$ interactions to generate a summary $S=f_{\theta}(X_{\text{enc}})$.
-We evaluate this summary on two disjoint subsets: (i) $X_{s}\subseteq X_{\text{enc}}$: past questions for assessing reconstruction ability; and (ii) $Y_{s}\subseteq Y$: future questions for evaluating predictive accuracy.
-The summary $S$ is passed to the frozen decoder $g$, which outputs predictions $\hat{X}_{s}=g(S;X_{s})$ and $\hat{Y}_{s}=g(S;Y_{s})$.
-Reward function.
-The reward for a summary $S$ is defined using a function $\Phi$ that balances multiple criteria:
-$\mathcal{R}(S;g)=\Phi\left(\texttt{Acc}(\hat{X}_{s},X_{s}),\;\texttt{Acc}(\hat{Y}_{s},Y_{s}),\;|S|,\;\Omega(S)\right),$ (1)
-where $\texttt{Acc}(\cdot)$ measures accuracy, $|S|$ is the summary length, and $\Omega(S)$ enforces structural constraints (e.g., inclusion of a Misconceptions section).
-The function $\Phi$ balances these components using hyperparameters or indicator functions to enforce constraints as needed.
-Optimization via GRPO.
-To optimize $f_{\theta}$, we employ Group Relative Policy Optimization (GRPO) shao2024deepseekmath .
-For each input $X$, the encoder generates $G$ candidate summaries $\{S_{1},\dots,S_{G}\}$, each evaluated using $\mathcal{R}(S_{i};g)$.
-We compute group-relative advantage and update parameters:
-$A(S_{i})=\frac{\mathcal{R}(S_{i};g)-\mu}{\sigma},\quad\nabla_{\theta}\mathbb{E}_{S\sim f_{\theta}(X)}[\mathcal{R}(S;g)]\approx\frac{1}{G}\sum_{i=1}^{G}A(S_{i})\cdot\nabla_{\theta}\log p_{\theta}(S_{i}|X),$ (2)
-where $\mu$ and $\sigma$ are the mean and standard deviation of rewards within the group. This approach encourages the encoder to produce summaries that maximize the reward function defined in Eq. 1.
-3.4 Steerability of the Estimated Knowledge States
-A key advantage of LBMs is the ability for humans to steer the model’s behavior. We identify three mechanisms for human-model interaction (detailed in Appendix):
-  • Prompt engineering the encoder. Directly shaping summary generation through input instructions.
-  • Steering via reward signals. Incorporating human preferences during encoder training through the $\Omega(S)$ term in Eq. 1.
-  • Augmenting with student-specific information.
-Educators can supplement the model with additional knowledge not present in observed data—either by augmenting encoder inputs or by editing the generated summary before decoding.
-This enables integration of recent classroom observations or specific misconceptions identified through in-person interactions.
-4 Related Work
-In this section, we review the literature related to our proposed approach, spanning traditional knowledge tracing, recent advances in LLM-based student modeling, and concept bottleneck models (CBMs), and text summarization models.
-Traditional Knowledge Tracing
-Traditional knowledge tracing (KT) methods have long been used to model student learning and predict future performance shen2024survey .
-Early works such as Bayesian Knowledge Tracing (BKT) corbett_knowledge_1994 model student mastery as a probabilistic process, while more recent methods like Deep Knowledge Tracing (DKT) piech_deep_2015 and Attentive Knowledge Tracing (AKT) ghosh2020context uses an attention-based architecture to learn contextual representation of questions and student knowledge states.
-Despite their strong predictive performance, these models represent a student’s knowledge state as an abstract, high-dimensional latent vector, which poses significant challenges in interpretability and actionable feedback for educators.
-LLM-Based Knowledge Tracing
-Recent studies have begun integrating Large Language Models (LLMs) into the KT framework.
-For example, li2024explainable demonstrated that LLMs are able to make sensible predictions about student responses when prompted with adequate information.
-Other works lee2024language ; kim2024beyond have studied how LLMs can help mitigate the cold-start problem compared to traditional KT approaches, while wang2025llm demonstrated state-of-the-art performances in KT by combining LLMs with sequence interaction models.
-However, these methods generally remain opaque: they either treat the LLM as a black-box, or rely on model-generated explanations that are susceptible to hallucination (bender2021dangers, ).
-Concept Bottleneck Models
-Concept Bottleneck Models (CBMs) koh2020concept improve interpretability through human-understandable concept activations as intermediates, with extensions exploring unsupervised concept learning oikarinen2023label , test-time interventions shin2023closer , and theoretical analyses of concept set design luyten2024theoretical .
-However, CBMs typically rely on finite predefined concept sets, limiting their applicability to complex tasks like knowledge tracing.
-Recently, yamaguchitoward introduced Explanation Bottleneck Models (XBMs), which use textual rationales as intermediates for vision classification, justifying a single known label per input.
-While our Language Bottleneck Models (LBMs) adopt this language bottleneck concept, they differ fundamentally.
-Unlike XBMs’ instance-specific, task-specific rationales, LBM summaries aim to capture an implicit knowledge states—as emphasized by our inverse problem formulation—and generalize to future, unknown questions.
-These requirements necessitate holistic, adaptable summaries rather than XBMs’ task-specific rationales, leading us to introduce Language Bottleneck Models (LBMs) as a distinct framework with broader applicability.
-Contrast with Text Summarization Models
-Recent text summarization models such as BART lewis2019bart , T5 raffel2020exploring , and PEGASUS zhang2020pegasus effectively produce concise summaries based on explicitly available textual content.
-However, these approaches differ fundamentally from knowledge tracing, where summaries must infer implicit student knowledge states not directly observable in the input.
-Standard summarization metrics (e.g., ROUGE, BLEU) rely on explicit reference summaries, making them unsuitable for evaluating latent knowledge inference tasks.
-In contrast, our Language Bottleneck Models generate textual summaries of the student’s implicit knowledge state, optimized for predictive accuracy on downstream questions rather than syntactic overlap with the observed input.
-5 Experiments
-We evaluate the effectiveness of our Language Bottleneck Models (LBM) in the context of Knowledge Tracing (KT), both with a controlled synthetic KT task involving basic arithmetic operations, and on real-world knowledge tracing datasets.
-5.1 Experimental Setup
-Datasets We evaluate on one synthetic and one real-world source.
-Synthetic arithmetic simulates learners answering addition, subtraction, multiplication, and division items; each student is assigned mastered skills, unmastered skills, and systematic misconceptions.
-The Eedi dataset wang2020diagnostic contains millions of question–answer pairs from UK secondary-school mathematics.
-More details about these datasets are provided in the Appendix.
-Models The main experiments use GPT-4o and GPT-4o-mini, together with open-source Qwen 2.5 (3B, 7B) and Gemma 3 (12B, 30B) checkpoints.
-Unless noted, we run the instruct variants of each open-source model; hyper-parameters and prompt templates are provided in the Appendix.
-Baselines We compare LBMs against (i) four standard knowledge-tracing methods—DKT piech_deep_2015 , DKVMN zhang2017dynamic , SAKT pandey2019self , and AKT ghosh2020context —implemented with the pyKT library liupykt2022 ; and (ii) direct prompting, where each LLM predicts answers from the full student history without a bottleneck.
-Training details for all baselines are given in Appendix.
-5.2 Results
-5.2.1 LBM vs Direct Prompting
-Table 1: Accuracy results for Direct and LBM methods. Bold indicates LBM accuracy is no more than 2% below the Direct baseline.
-synthetic eedi Direct LBM $\Delta$ Direct LBM $\Delta$ Qwen2.5-3B 0.525 0.627 0.102 0.347 0.372 0.025 Qwen2.5-7B 0.655 0.646 -0.009 0.647 0.583 -0.065 gemma-3-12b 0.790 0.791 0.001 0.580 0.625 0.045 gemma-3-27b 0.785 0.794 0.009 0.670 0.650 -0.020 gpt-4o-mini 0.795 0.782 -0.013 0.655 0.645 -0.010 gpt-4o 0.840 0.825 -0.015 0.703 0.682 -0.020
-Figure 4: Evolution of accuracy with bottleneck length.
-Table 1 compares the accuracy of LBM models to direct prompting variants across different LLMs.
-Despite the information bottleneck enforced by LBMs, they remain within $2\%$ of the accuracy obtained via direct prompting.
-Figure 4 demonstrates that this gap decreases for longer bottlenecks, highlighting a trade-off between conciseness and predictive accuracy.
-Encoder/decoder variants
-We evaluate LBMs with different models as encoders and decoders (results shown in the Appendix).
-Using the strongest model (gpt-4o) as encoder with other models from Table 1 as decoders yields accuracies $5-10\%$ higher than when the stronger model is used as the decoder.
-This confirms our hypothesis that extracting relevant information for the summary (encoding) is more challenging than predicting future answers given a summary (decoding).
-5.2.2 LBM vs Knowledge Tracing Methods
-Figure 5 compares the performance of LBM models to traditional Knowledge Tracing methods on the synthetic dataset.
-LBMs with strong LLMs achieve comparable accuracy to KT methods with significantly less training data.
-Since KT methods rely on statistical patterns, they require substantially more observations before reaching strong predictive power, while LBMs demonstrate effective generalization with minimal examples.
-Figure 5: Accuracy of LBM vs Knowledge Tracing models on the Synthetic dataset.
-The x axis represents the total number of question seen by the model, ie the product of number of questions observed per student $T$ and the number of students available for training $N$ (note that N=1 for LBM methods).
-The y axis is the average accuracy 200 test student and 4 new questions to predict per student.
-The pre-trained models indicated for LBM are used as both Encoder and Decoder, with chain-of-thought prompting.
-Error bars report the standard error calculated across 10 runs for N=10 students and 3 runs otherwise.
-5.2.3 Training LBM Encoders
-As discussed in Section X, downstream accuracy serves as a reward signal to train encoders to produce increasingly accurate summaries.
-We demonstrate this by training a Gemma3-12B encoder with GRPO alongside a frozen Gemma3-27B decoder.
-Figure 6 shows the encoder progressively improving summary quality, quickly outperforming gpt-4o.
-The figure illustrates this progression through an example student who mastered all constructs except subtraction and fails any multiplication involving 9.
-The initial summary contains inaccuracies and misses this systematic misconception, while the final summary successfully captures the student’s complete knowledge state.
-Figure 6: Training the encoder with the decoding accuracy on the synthetic dataset.
-Evolution of the test accuracy as a Gemma3-12B encoder is trained as described in Section 3.3, with a fixed Gemma3-27B decoder.
-Trained on $800$ training students and tested on $200$ students, with $|\mathcal{X}|=50$ questions per input trajectory and $|\mathcal{Y}|=20$ questions to predict per student.
-The bottom row show the evolution of the generated summary over the course of training for an example student.
-Text is colored green (exact), orange (approximate), or red (false) based on ground-truth.
-Accuracy gains from training are stronger on harder cases
-When stratifying students by difficulty (measured by number of misconceptions), we find training yields larger accuracy improvements for complex cases (see table in the Appendix).
-The model shows 10-11% increased accuracy relative to gpt-4o for students with 2 or 3+ misconceptions, compared to just 5-7% for students with 0-1 misconceptions.
+Bottom right notations correspond to components in Figure 1.
+A.2 Full results
+Table A.2 compares the zero-shot performance of LBMs to all the baselines across all datasets.
+Note: The surprisingly low accuracy of some LLMs on EEDI and XES3G5M likely reflects the strong class imbalance in these datasets (68% correct in EEDI; 85% in XES3G5M; cf.
+Table 4) and the fact that both LBMs and Direct models are evaluated zero-shot, preventing them from adjusting to this imbalance.
+Consequently, models whose predictions skew toward the minority class (“incorrect”) can score below random accuracy.
+The high AUCs for some of these models indicate that they still capture some meaningful predictive signal.
+Table A1: Results across all datasets and methods.
+We report the mean and standard deviation across N=10 runs for KT and CD models and N=3 runs for LLM-based models (Direct, LBM).
+XES3G5M-E and XES3G5M-C denote the English and Chinese versions of the XES3G5M dataset, respectively; note that since the KT and CD models do not use the question content, their results are shared across both versions of the dataset.
+"Zero-shot" methods (LLM Direct; most LBMs) are run zero-shot on a test set; "Trained" methods (KT and CD models, trained LBM) are trained on a train set and evaluated on a test set.
+For the trained LBM (last row), "E" and "D" stand for Encoder and Decoder respectively.
+"*" indicates the part of the model that is trained.
+AUC is unavailable for closed-source LLM models as they require access to the model’s output logits.
+Results for QIKT on Synthetic and MIRT on EEDI are omitted due to implementation issues.
+Synthetic EEDI (Filt.) XES3G5M-E (Filt.) XES3G5M-C (Filt.) Model Type Model Name ACC AUC ACC AUC ACC AUC ACC AUC KT Models (Trained) akt .87±.01 .87±.01 .72±.02 .66±.02 .87±.01 .68±.03 .87±.01 .68±.03 deep irt .82±.02 .82±.01 .68±.02 .61±.02 .85±.01 .64±.03 .85±.01 .64±.03 dkt .82±.02 .81±.02 .66±.02 .57±.02 .85±.01 .64±.03 .85±.01 .64±.03 dkvmn .82±.01 .82±.01 .67±.03 .58±.02 .85±.01 .67±.03 .85±.01 .67±.03 gkt .71±.07 .67±.08 .63±.05 .58±.03 .86±.01 .60±.06 .86±.01 .60±.06 qikt .68±.00 .57±.00 .82±.00 .53±.00 .82±.00 .53±.00 saint .66±.02 .65±.02 .70±.02 .59±.05 .87±.01 .66±.03 .87±.01 .66±.03 sakt .82±.02 .81±.02 .67±.01 .56±.02 .85±.01 .60±.02 .85±.01 .60±.02 simplekt .86±.01 .85±.01 .72±.02 .63±.03 .86±.01 .68±.02 .86±.01 .68±.02 CD Models (Trained) DINA .58±.01 .63±.01 .51±.00 .55±.01 .50±.01 .55±.01 .50±.01 .55±.01 IRT .61±.01 .64±.01 .67±.00 .62±.01 .83±.00 .68±.01 .83±.00 .68±.01 KaNCD .90±.01 .96±.01 .65±.01 .64±.01 .85±.00 .81±.00 .85±.00 .81±.00 MIRT .63±.01 .68±.01 .71±.01 .66±.01 .71±.01 .66±.01 NCDM .90±.01 .96±.00 .66±.01 .67±.00 .85±.00 .80±.00 .85±.00 .80±.00 LLM Direct (Zero-shot) Qwen2.5-3B .56±.00 .79±.00 .35±.00 .54±.00 .19±.00 .56±.00 .19±.00 .55±.00 Qwen2.5-7B .64±.00 .73±.00 .65±.00 .62±.00 .76±.00 .63±.00 .75±.00 .63±.00 gemma-3-12b .63±.00 .96±.00 .58±.00 .71±.00 .30±.00 .71±.00 .29±.00 .70±.00 gemma-3-27b .62±.00 .94±.00 .67±.00 .76±.00 .33±.00 .78±.00 .32±.00 .79±.00 gpt-4o .85±.01 .66±.00 .82±.00 .79±.01 gpt-4o-mini .81±.01 .67±.01 .74±.08 .75±.07 gpt-5 .87±.00 .71±.02 .81±.01 .79±.01 LBM (Zero-shot) Qwen2.5-3B .61±.01 .65±.01 .38±.00 .55±.01 .27±.01 .58±.01 .30±.01 .63±.02 Qwen2.5-7B .65±.01 .69±.01 .58±.01 .55±.01 .75±.01 .63±.00 .74±.01 .62±.02 gemma-3-12b .79±.00 .85±.00 .62±.01 .64±.02 .63±.01 .67±.01 .57±.02 .65±.01 gemma-3-27b .78±.01 .85±.01 .65±.01 .67±.01 .68±.02 .70±.01 .68±.01 .70±.02 gpt-4o .80±.01 .66±.02 .80±.01 .78±.02 gpt-4o-mini .78±.01 .61±.02 .70±.08 .70±.07 gpt-5 .87±.01 .68±.01 .76±.01 .76±.01 LBM (Trained) E: gpt-5, D: gemma-3-12b* .70±.02 .73±.01 .83±.01 .76±.01
+A.3 Training LBM Encoders
+A.3.1 Difficulty stratification
+To investigate whether the accuracy gains seen during LBM training (Figure 5) vary across students in the dataset, we stratify students according to how many misconception they hold: 0, 1, 2 or 3+.
+Since each misconception represent additional "irregularities" in the student’s response pattern beyond simple mastery of constructs, this effectively stratifies different complexity levels across student knowledge states.
+Table A2 shows the change in accuracy relative to the GPT-4o baseline across difficulty levels.
+Table A2: Relative difference in accuracy between trained Gemma3-12B and the GPT-4o baseline, with students grouped by number of misconceptions.
+Accuracy (mean ± std) Relative Diff. (%)
+# of Misconceptions Baseline Pre-training Post-training Pre Post $\Delta$
+0   ($N=22$) 0.98 ± 0.05 0.92 ± 0.12 0.99 ± 0.05 -5.80 0.90 6.70
+1   ($N=17$) 0.89 ± 0.09 0.87 ± 0.15 0.91 ± 0.09 -2.00 3.00 5.00
+2   ($N=23$) 0.81 ± 0.16 0.80 ± 0.12 0.89 ± 0.08 -0.80 10.20 11.00
+3   ($N=135$) 0.77 ± 0.12 0.74 ± 0.14 0.82 ± 0.11 -4.80 5.70 10.50
+The model shows $10-11\%$ increased accuracy relative to GPT-4o for students with 2 or 3+ misconceptions, compared to just $5-7\%$ for students with 0-1 misconceptions.
 This suggests the encoder becomes particularly better at handling complex knowledge states.
-5.2.4 Steering LBM behavior
-We demonstrate multiple steering strategies described in Section 3.4 (details in Appendix).
-We find that providing explicit misconception information during encoder training produces significantly stronger learning effects than adding the same information at the decoder stage.
-This suggests the encoder leverages this additional context to better interpret patterns in student responses.
-We also show the encoder can be steered during training to explicitly mention misconceptions through reward signals, and illustrate how supplementing the bottleneck with information not present in the input can significantly improve decoder accuracy.
-6 Discussion
-Core contributions
-Framing Knowledge Tracing as an inverse problem shifts focus from fitting statistical regularities to understanding the underlying reasons for student responses.
-Language Bottleneck Models operationalize this view by compressing interaction histories into explicit natural-language summaries that maintain high predictive power while being both interpretable and steerable by domain experts.
-Why can’t we just prompt GPT-4o directly?
-The split encoder-decoder architecture of LBMs offers three key advantages over direct LLM prompting: it creates a global student model with a single latent summary shared across all predictions rather than isolated per-question reasoning; it ensures faithful summaries through a closed-loop decoding objective that penalizes non-predictive summaries; and it provides an explicit interface layer that teachers can read, steer and intervene on.
-Wider applicability
-LBMs extend beyond education to any task requiring compact, human-readable summaries with predictive power.
-The minimal ingredients needed are: (1) a sequence of observations about an entity, (2) a need to predict future behaviors of that entity, and (3) value in having interpretable representations.
-For example, clinical decision support could distill patient data into textual state descriptions that forecast outcomes while remaining auditable; preventive maintenance could compress sensor logs into explanations predicting machine failure; customer success teams could summarize interaction histories to forecast churn.
-Limitations
-LBMs face several constraints: context length limitations imposed by underlying LLMs, which could be addressed through iterative summarization; the requirement for textual question content rather than just identifiers, limiting applicability to some KT datasets; and the substantial computational cost of running LLMs.
-We discuss limitations extensively in the Appendix.
-Future work
-Several promising extensions include: implementing iterative encoding to process longer inputs through sequential chunking; developing active sensing capabilities to select questions that maximize information about students’ knowledge states; adapting LBMs for evolving knowledge states to capture learning and forgetting over time; expanding beyond question-answer inputs to leverage richer data sources; and combining with pedagogical-alignment techniques like LearnLM (team2024learnlm, ).
-We discuss these points in more details in the Appendix.
-By bridging the gap between predictive power and human interpretability, Language Bottleneck Models represent a significant step toward AI systems that can genuinely augment human expertise in education and beyond.
-Appendix A Introduction to Knowledge Tracing
-In this section, we provide a more detailed introduction to Knowledge Tracing. For recent comprehensive reviews of the field, see [1, 25].
-Knowledge Tracing (KT) is the task of modeling students’ knowledge states and predicting their performance on future questions based on their historical interactions with educational content.
-The goal is to dynamically track the acquisition, retention, and application of knowledge as students engage with learning materials.
-A.1 Formal Problem Definition
-Knowledge Tracing (KT) is fundamentally a sequence modeling problem where the objective is to infer a student’s latent knowledge state based on observed interactions.
-Formally, let $\mathcal{X}=\{(q_{1},a_{1}),(q_{2},a_{2}),\ldots,(q_{t},a_{t})\}$ denote a sequence of question-answer pairs, where $q_{i}\in\mathcal{Q}$ represents a question from the question space $\mathcal{Q}$, and $a_{i}\in\mathcal{A}$ represents the student’s answer from the answer space $\mathcal{A}$.
-The goal of KT is to learn a function that predicts the probability of correctly answering a future question $q_{t+1}$ given the historical interaction sequence: $P(a_{t+1}|q_{t+1},\mathcal{X})$.
-In practice, each interaction $(q_{i},a_{i})$ in the dataset typically contains richer information.
-Questions $q_{i}$ often include metadata about the underlying knowledge constructs they assess (e.g., specific skills, concepts, or learning objectives).
-For example, a mathematics question might be tagged with constructs like "quadratic equations" or "algebraic factoring." Similarly, observed answers $a_{i}$ can vary in granularity—from binary correctness indicators to specific answer choices in multiple-choice questions, or complete textual/code responses for open-ended questions.
-Many KT datasets also capture additional contextual information such as response time or self-reported confidence scores, which can be leveraged by KT models.
-The prediction target can also extend beyond binary correctness.
-Models may predict specific answer choices a student will select (see for example Task 2 of Eedi’s NeurIPS 2020 Education Challenge [30]), enabling more nuanced understanding of misconceptions.
-In domains like programming education, models might predict the exact code a student will write [14], capturing finer-grained knowledge states.
-A.2 Knowledge Tracing Paradigms
-Bayesian Knowledge Tracing (BKT).
-Introduced by [5], BKT models student knowledge as a set of binary latent variables corresponding to specific skills or concepts.
-For each skill, a student is either in a "known" or "unknown" state.
-BKT employs a Hidden Markov Model to update these probabilities based on observed student responses, accounting for learning, guessing, and slipping in its probabilistic framework.
-This approach provides interpretable outputs, with each concept having a probability of mastery, but is limited in modeling the complex interdependencies between skills.
-Deep Learning Approaches.
-Deep Knowledge Tracing (DKT) [20] marked a paradigm shift by using recurrent neural networks to model student knowledge as continuous latent representations.
-Unlike BKT’s discrete skill probabilities, DKT captures complex patterns and skill relationships through high-dimensional embeddings, significantly improving prediction accuracy.
-Subsequent models have enhanced this approach, including: Dynamic Key-Value Memory Networks (DKVMN) [34] which incorporate memory components for better concept tracking; Self-Attentive Knowledge Tracing (SAKT) [19] and Attentive Knowledge Tracing (AKT) [6] which leverage attention mechanisms to capture long-range dependencies; and Graph-based Knowledge Tracing models [17, 33] which incorporate structural information into knowledge representations.
-A.3 Current Challenges
-Interpretability vs. Accuracy Trade-off.
-Deep learning approaches typically achieve higher predictive accuracy but at the cost of interpretability.
-Their latent representations are opaque high-dimensional vectors that provide limited insight into specific concepts a student has mastered or misconceptions they hold.
-This limits the usability of these models in practice.
-Why do we need interpretability?
-At first glance, predictive accuracy alone might seem sufficient for KT—an accurate model could, for example, analyze student responses across many questions to identify overall strengths and weaknesses.
-However, even highly accurate models can yield misleading or fragile predictions if the underlying reasoning is opaque.
-For instance, [22] showed that predictive accuracy alone might obscure models relying on superficial patterns like repeated questions, limiting practical usability.
-Thus, without interpretability, accurate models may fail to meaningfully support educators in diagnosing and intervening in students’ learning processes.
-Cold-start Problem.
-KT systems struggle with new students or questions where historical data is limited or non-existent: for any new question added to the dataset, traditional models require the observation of answers from many students to learn an accurate latent representation for this question.
-This "cold-start" problem poses significant challenges for the practical deployment of KT systems, particularly in dynamic educational environments where new content and students are regularly introduced.
-Appendix B Extended Results
-B.1 Encoder/Decoder Variants (Section 5.2.1)
-Table B1 shows the accuracy of different combinations of encoder-decoder models on the Synthetic dataset.
+A.3.2 Evolution of misconception detection and false positive
+Figure A2 shows the evolution of the misconception detection rate and misconception false positive counts over the training of the encoder from §4.2.
+Figure A2: Evolution of misconception detection and false positive over encoder training.
+A.4 Steering LBM behavior
+A.4.1 Augmenting with student-specific information to assist LBM training
+Finally, we demonstrate how providing additional information to the LBM can assist with the training process.
+On the Synthetic dataset, we train two identical LBMs (Gemma-12B trainable encoder, Gemma-27B frozen decoder) while providing 2 misconceptions either to the encoder as part of the input data, or to the decoder as part of the bottleneck.
+The models are trained for one epoch on $800$ students.
+We then evaluate the resulting models without any additional information provided.
+Table A3 shows the accuracy before/after training for both models.
+The model where additional information was provided to the encoder during training reaches $84\%$ accuracy after training compared to $80\%$ when it is provided to the decoder.
+This suggest that the additional information facilitates training of the encoder.
+Table A3: Result of training on the Synthetic dataset when providing misconception information about each student during training, either in the encoder input or in the bottleneck.
+Before training After 1 epoch training
+Information in X 0.802 ± 0.001 0.840 ± 0.004
+Information in S 0.794 ± 0.002 0.799 ± 0.006
+A.4.2 Steering via reward signals
+To demonstrate the possibility to steer LBMs’ behavior via the reward signal, we consider an example use-case where a teacher would like the model to pay particular attention to potential misconceptions held by the student.
+Following the reward-shaping framework of Section 3.3, we augment the training objective with an additional term that explicitly encourages the model to surface misconceptions.
+Concretely, we set $\Omega(S)=\mathbbm{1}\!\bigl[\text{``misconception''}\in\tilde{\mathcal{S}}\bigr$], where $\tilde{\mathcal{S}}$ is the textual bottleneck emitted by the LBM for student state $S$.
+This binary reward is added to the accuracy term and optimized with GRPO.
+Figure A3 confirms that the policy quickly internalizes this incentive: after only a handful of training steps, the proportion of summaries that explicitly mention a misconception goes from less than 80% to >95%, demonstrating that the reward function provides an effective lever for shaping higher-level pedagogical behavior.
+Figure A3: Presence of the word "misconception" in the LBM’s summaries during training of the encoder model steered towards mentioning student misconceptions via the reward signal.
+A.4.3 Augmenting LBMs with student-specific information to complement input data
+To demonstrate how an LBM can be actively steered with additional information we run the following ablation on the Synthetic dataset: 1.
+Each student trajectory originally probes four constructs.
+For a given trajectory, we remove every question linked to one construct leaving the input data intentionally incomplete.
+2.
+We then run the input data through the encoder and inject a single teacher-supplied sentence describing the student’s mastery of the missing construct directly into the model’s bottleneck representation (e.g., "The student has mastered addition except in the event of misconceptions").
+We repeat this procedure four times—once for each construct—running the LBM both with and without the additional sentence, and report the results Table A4.
+Naturally, the models complemented with the additional information of the student’s mastery of the construct missing in the input data outperform models provided only with the incomplete data.
+This small experiment illustrates a key advantage of LBMs: because the LBM compresses evidence into a text-based summary, it can be complemented with additional information absent from the input data.
+Table A4: Accuracy of a gpt-4o-based LBM on the Synthetic dataset while removing one construct from the input data, with and without providing information about the missing construct in the bottleneck.
+Each run is repeated across all four constructs; mean and standard deviation are reported.
+Without additional bottleneck information With additional bottleneck information
+0.749 ± 0.007 0.791 ± 0.013
+A.5 Ablation Experiments
+A.5.1 Encoder/Decoder Variants
+Table A5 shows the accuracy of different combinations of encoder-decoder models on the Synthetic dataset.
 The top row shows the result of using the strongest model evaluated (gpt-4o) as both encoder and decoder.
@@
 This implies that the task of accurately capturing knowledge state information in the bottleneck is harder than predicting answers to future questions when provided with a knowledge state summary.
-Table B1: Performance of different LBM encoder-decoder combinations on the Synthetic dataset.
+Table A5: Performance of different LBM encoder-decoder combinations on the Synthetic dataset.
 Encoder Decoder Accuracy
@@
 decoder Qwen/Qwen2.5-3B-Instruct gpt-4o 0.649
-B.2 Traditional KT methods on the Eedi dataset
-We run the four benchmark KT methods on the Eedi dataset with the same experimental setup as for Table 1 in the main paper (see Section D.2.2 in this appendix for dataset details).
-Table B2 shows the resulting accuracies, alongside the accuracy found with Direct and LBM models for different LLMs, reproduced from Table 1.
-The best KT model (AKT) slightly outperform the best LBM and Direct models with gpt-4o, while other KT models reach similar accuracies to LBM and Direct models with gpt models and Gemma3-27B.
-Table B2: Average accuracy on the Eedi dataset for different model types.
-For each student, the accuracy is computed on the last four questions of the trajectory.
-KT models show mean $\pm$ SEM across 10 runs.
-Due to computational constraints, Direct and LBM results are reported from single runs.
-Model Accuracy
-KT Models
-DKT $0.647\pm 0.005$
-DKVMN $0.669\pm 0.004$
-SAKT $0.680\pm 0.007$
-AKT $0.720\pm 0.008$
-Direct
-Qwen2.5-3B 0.347
-Qwen2.5-7B 0.647
-Gemma3-12B 0.580
-Gemma3-27B 0.670
-GPT-4o-mini 0.655
-GPT-4o 0.703
-LBM
-Qwen2.5-3B 0.372
-Qwen2.5-7B 0.583
-Gemma3-12B 0.625
-Gemma3-27B 0.650
-GPT-4o-mini 0.645
-GPT-4o 0.682
-B.3 Training LBM Encoders - Difficulty stratification (Section 5.2.3)
-To investigate whether the accuracy gains seen during LBM training (cf Figure 6 in the main paper) vary across students in the dataset, we stratify students according to how many misconception they hold: 0, 1, 2 or 3+.
-Since each misconception represent additional "irregularities" in the student’s response pattern beyond simple mastery of constructs, this effectively stratifies different difficulty levels across students.
-Table B3 shows the change in accuracy relative to the gpt-4o baseline across difficulty levels.
-Table B3: Relative difference in accuracy between trained Gemma3-12B and the GPT-4o baseline, with students grouped by number of misconceptions.
-Accuracy (mean ± std) Relative Diff. (%)
-# of Misconceptions Baseline Pre-training Post-training Pre Post $\Delta$
-0   ($N=22$) 0.98 ± 0.05 0.92 ± 0.12 0.99 ± 0.05 -5.80 0.90 6.70
-1   ($N=17$) 0.89 ± 0.09 0.87 ± 0.15 0.91 ± 0.09 -2.00 3.00 5.00
-2   ($N=23$) 0.81 ± 0.16 0.80 ± 0.12 0.89 ± 0.08 -0.80 10.20 11.00
-3   ($N=135$) 0.77 ± 0.12 0.74 ± 0.14 0.82 ± 0.11 -4.80 5.70 10.50
-B.4 Steering LBM behavior
-B.4.1 Steering via reward signals
-To demonstrate the possibility to steer LBMs’ behavior via the reward signal, we consider an example use-case where a teacher would like the model to pay particular attention to potential misconceptions held by the student.
-Following the reward-shaping framework of Section 3.3, we augment the training objective with an additional term that explicitly encourages the model to surface misconceptions.
-Concretely, we set $\Omega(S)=\mathbbm{1}\!\bigl{[}\text{``misconception''}\in\tilde{\mathcal{S}}\bigr{]}$, where $\tilde{\mathcal{S}}$ is the textual bottleneck emitted by the LBM for student state $S$.
-This binary reward is added to the accuracy term and optimized with GRPO.
-Figure B1 confirms that the policy quickly internalizes this incentive: after only a handful of training steps, the proportion of bottlenecks that explicitly mention a misconception goes from less than 80% to >95%, demonstrating that the reward function provides an effective lever for shaping higher-level pedagogical behavior.
-Figure B1: Presence of the word "misconception" in the LBM’s bottlenecks during training of the encoder model steered towards mentioning student misconceptions via the reward signal.
-B.4.2 Augmenting LBMs with student-specific information to complement input data
-To demonstrate how an LBM can be actively steered with additional information we run the following ablation on the Synthetic dataset: 1.
-Each student trajectory originally probes four constructs.
-For a given trajectory, we remove every question linked to one construct leaving the input data intentionally incomplete.
-2.
-We then run the input data through the encoder and inject a single teacher-supplied sentence describing the student’s mastery of the missing construct directly into the model’s bottleneck representation (e.g., "The student has mastered addition except in the event of misconceptions.).
-We repeat this procedure four times—once for each construct—running the LBM both with and without the additional sentence, and report the results Table B4.
-Naturally, the models complemented with the additional information of the student’s mastery of the construct missing in the input data outperform models provided only with the incomplete data.
-This small experiment illustrates a key advantage of LBMs: because the LBM compresses evidence into a text-based bottleneck, it can be complemented with additional information absent from the input data.
-Table B4: Accuracy of a gpt-4o-based LBM on the Synthetic dataset while removing one construct from the input data, with and without providing information about the missing construct in the bottleneck.
-Each run is repeated across all four constructs; mean and standard deviation are reported.
-Without additional bottleneck information With additional bottleneck information
-0.749 ± 0.007 0.791 ± 0.013
-B.4.3 Augmenting with student-specific information to assist LBM training
-Finally, we demonstrate how providing additional information to the LBM can assist with the training process.
-We train two identical LBMs while providing 2 misconceptions either to the encoder as part of the input data, or to the decoder as part of the bottleneck.
-The models are trained for one epoch on $800$ students.
-We then evaluate the resulting models without any additional information provided.
-Table B5 shows the accuracy before/after training for both models.
-The model where additional information was provided to the encoder reaches $84\%$ accuracy after one training compared to $80\%$ when it is provided to the decoder.
-This suggest that the additional information facilitates training of the encoder.
-Table B5: Relative difference in accuracy between trained Gemma3-12B and the gpt-4o baseline, with students grouped by number of misconceptions.
-Before training After 1 epoch training
-Information in X 0.802 ± 0.001 0.840 ± 0.004
-Information in S 0.794 ± 0.002 0.799 ± 0.006
-B.5 Example bottlenecks on the Eedi dataset
-Figure B2 shows example bottlenecks produced by a gpt-4o-based LBM for a student in the Eedi dataset.
+A.5.2 Importance of Knowledge Concepts in the input
+Table LABEL:app:tab:ablation compares the accuracy of different LBMs with vs without knowledge conception information in the input prompt.
+The performance generally remains similar to the KC version, and it even increases for most models on the EEDI dataset.
+This demonstrates that LBMs do not fundamentally require KC information.
+Table A6: Comparison of LBM performance with vs without knowledge concept (KC) information in the input. Bold indicates a difference in accuracy no more than $3\%$ below the full input.
+Synthetic EEDI (Filtered)
+w/ KC w/o KC $\Delta$ w/ KC w/o KC $\Delta$
+Qwen2.5-3B .61 ± .01 .62 ± .05 +.00 .38 ± .00 .38 ± .02 +.01
+Qwen2.5-7B .65 ± .01 .66 ± .02 +.01 .58 ± .01 .58 ± .01 +.00
+gemma-3-12b .79 ± .00 .73 ± .00 -.06 .62 ± .01 .60 ± .02 -.02
+gemma-3-27b .78 ± .01 .75 ± .01 -.03 .65 ± .01 .69 ± .03 +.03
+gpt-4o-mini .78 ± .01 .76 ± .00 -.01 .61 ± .02 .65 ± .03 +.04
+gpt-4o .80 ± .01 .76 ± .01 -.04 .66 ± .02 .68 ± .01 +.02
+A.6 Example bottlenecks on the Eedi dataset
+Figure A4 shows example bottlenecks produced by a gpt-4o-based LBM for a student in the Eedi dataset.
 The input data is composed of 30 questions from various constructs, and the decoder predicts 4 test questions.
@@
 A longer bottleneck of 256 or 512 tokens allows for more nuance and details in the bottleneck, leading to the decoder correctly predicting all four test questions for this student.
-Figure B2: Example bottlenecks produced by an LBM with gpt-4o as both encoder and decoder for a student in the Eedi dataset, with varying bottleneck lengths.
+Figure A4: Example bottlenecks produced by an LBM with gpt-4o as both encoder and decoder for a student in the Eedi dataset, with varying bottleneck lengths.
 The input data provided to the encoder is composed of 30 questions across various constructs, and the decoder predicts 4 test questions using the bottleneck.
-Appendix C Extended Discussion
-C.1 Extended Discussion
-Knowledge Tracing as an Inverse Problem
-The reframing of Knowledge Tracing as an inverse problem represents a fundamental shift in how we conceptualize student modeling.
-Traditional KT methods adopt a forward modeling approach, where the goal is to learn a mapping from observed interactions to future performance predictions.
-While effective for prediction, this treats the student’s knowledge state as an opaque latent variable that serves primarily as an intermediate computational step.
-In contrast, our inverse problem formulation explicitly seeks to recover the underlying knowledge state that generates observed behaviors, making interpretability the primary objective rather than an afterthought.
-This aligns with educational assessment principles, where the main objective is to understand what a student knows, rather than merely predicting future behavior.
-Parallel with Kolmogorov Complexity
-Our inverse problem formulation of Knowledge Tracing draws a natural parallel with algorithmic information theory.
-The Kolmogorov complexity $K(x)$ of a string $x$ is defined as the length of the shortest program that outputs $x$ when run on a universal Turing machine.
-Traditional KT methods focus on learning statistical regularities between questions and responses, without looking for parsimonious explanations for observed behavior.
-In contrast, LBMs explicitly search for minimal natural language descriptions that can both reconstruct past interactions and predict future responses.
-The knowledge state summary $\mathcal{S}$ is akin to the minimal program, and the decoder LLM functions is akin to the universal interpreter that unpacks $\mathcal{S}$ to generate the observed question-answer patterns.
-This connection suggests that effective knowledge state summaries should capture the algorithmic essence of student behavior—the underlying "program" of knowledge and misconceptions that generates observable responses—rather than merely fitting surface-level patterns.
-While true Kolmogorov complexity is uncomputable, LBMs approximate this ideal through the natural language bottleneck constraint, encouraging summaries that balance compression with predictive power.
-Computational Architecture and Design Choices
-The encoder-decoder architecture reflects a deliberate separation of concerns: extraction versus interpretation of knowledge states.
-Our experiments demonstrate that these two tasks have different intrinsic complexity, with encoding (summary generation) being more challenging than decoding (prediction from summaries).
-This finding has practical implications for model selection and computational resource allocation.
-The use of GRPO for encoder training represents a novel application of reinforcement learning to interpretable AI, where the reward signal directly measures the downstream utility of explanations rather than their surface-level quality.
-C.2 Extended Limitations
-Context Length Constraints
-A practical limitation of LBMs is the context length restrictions of current LLMs.
-Eventhough modern models can handle thousands of tokens, comprehensive student histories in real educational settings can easily exceed these limits.
-As student trajectories grow more complex and include more information, the context required for the encoder might go beyond what current LLMs are capable of.
-A natural solution could be an iterative encoding process, where the encoder iterates over the text bottleneck while going over the input data by windows.
-Computational Cost and Resource Requirements
-Since LBMs involve two LLMs as the encoder and decoder, they are typically more computationally intensive to run than traditional KT methods.
-Each inference call with LBMs requires two LLM calls (encoding and decoding) with extensive context windows.
-Training LBMs with GRPO is particularly costly, as it requires iteratively generating and evaluating multiple candidate summaries per training example.
-This cost structure may limit the practical deployment of LBMs to high-stakes educational contexts where the interpretability benefits justify the computational expense.
-Dependence on Textual Question Content
-Many existing KT datasets provide only question identifiers rather than full question text.
-This limits the direct applicability of LBMs, as they require the questions content in order to generate meaningful summaries about student knowledge.
-While question content is increasingly available in modern educational platforms, this dependency creates a barrier to applying LBMs to historical datasets or systems that rely primarily on item response theory frameworks.
-Quasi-Knowledge State Assumption
-The quasi-static knowledge state assumption underlying our inverse problem formulation, while reasonable for short diagnostic sessions, does not cover longer time horizons.
-Real learning involves continuous knowledge acquisition, forgetting, and misconception evolution that our current framework cannot capture.
-This limitation restricts current LBMs to diagnostic rather than contexts of assessment throughout the learning process.
-The assumption also fails to account for contextual factors (fatigue, motivation, external stressors) that can significantly impact student performance within even short sessions.
-C.3 Extended Future Work
-Iterative Encoding Strategies
-A natural extension to address context length limitations involves iteratively applying the encoder while chunking input data.
-At each step, the encoder would process a portion of the interaction history $X$ along with one or more previously generated bottleneck summaries, creating an updated summary that incorporates new information from the latest chunk.
-This approach could maintain both detailed recent context and compressed historical patterns, enabling LBMs to handle arbitrarily long student trajectories through sequential refinement of knowledge state representations.
-Active Learning and Question Selection
-Building on iterative encoding, active learning strategies could provide encoders with the most informative input questions.
-Rather than processing all available interactions, the system could strategically select questions that maximize information gain about uncertain aspects of student knowledge.
-This could leverage uncertainty quantification methods or use the encoder LLM itself to recommend the most diagnostically valuable questions.
-Such active sensing would improve both efficiency and diagnostic power by focusing computational resources on the most insightful student responses.
-Dynamic Knowledge State Modeling
-Extending LBMs to handle evolving knowledge states represents a significant next step necessary to account for progressive learning and forgetting effects.
-A natural relaxation of the static knowledge state assumption involves "piecewise-constant" knowledge states that can evolve between question sessions but remain static within sessions.
-This extension poses exciting challenges in developing training objectives that balance within-session consistency with between-session learning dynamics, potentially requiring new approaches to temporal modeling in natural language representations.
-Multimodal Input Integration
-Question-answer pairs represent a relatively narrow information source for inferring student knowledge states.
-Richer data sources such as student-tutor interactions, self-reported explanations for answers, timing patterns, or hint usage could provide deeper insights into student understanding.
-While LBMs naturally extend to text-based inputs, future work should investigate optimal strategies for combining these varied data sources and evaluate their relative contributions to knowledge state inference accuracy and interpretability.
-Pedagogical Alignment
-Recent work like LearnLM [28] demonstrates the potential for making LLMs more pedagogically aligned.
-Incorporating similar pedagogical principles into LBM components could help encoders better interpret educational interactions and generate summaries that align with expert teaching practices.
-This might involve training on educator-annotated examples, incorporating educational taxonomies into summary structure, or using reward functions that emphasize pedagogically relevant aspects of student knowledge states.
-Such alignment could bridge the gap between computational convenience and educational validity.
-Appendix D Experimental Details
-D.1 Models
-LLM models
+Appendix B Experimental Details
+B.1 Language Bottleneck Models
+LLM backbones
 We evaluate the following closed- and open-source models:
-  • GPT-4o and GPT-4o-mini[2];
-  • Qwen2.5-3B and Qwen2.5-7B[29];
-  • Gemma3-12B and Gemma3-27B[27].
+  • GPT-5, GPT-4o and GPT-4o-mini(Achiam et al., 2023);
+  • Qwen2.5-3B and Qwen2.5-7B(Team, 2024);
+  • Gemma3-12B and Gemma3-27B(Team, 2025).
 For open-source models we extract the activation logits of the “Yes” and “No” tokens and return the higher value.
+The logits for the “Yes” and “No” tokens are used to compute the AUC.
 For closed-source models we prompt for a “Yes” or “No” answer and parse the text output.
+We run GPT models via the OpenAI API, with the following snapshots: GPT-5: 2025-08-07; GPT-4o: 2024-08-06; GPT-4o-mini: 2024-07-18.
+Reward function
+For training the encoder in Section 5, we set the reward function to the decoder accuracy across $|\mathcal{Y}|=20$ questions:
+$R(\tilde{\mathcal{S}};g)=\texttt{Acc}\big(\tilde{\mathcal{Y}},\mathcal{Y}\big)$
+The RL steering experiment section A.4.2 only rewards the presence of the word "misconception" in the bottleneck:
+$R(\tilde{\mathcal{S}};g)=\Omega(\tilde{\mathcal{S}})=\mathbbm{1}[\text{'misconception'}\in\tilde{\mathcal{S}}$]
+B.2 Training the decoder
+To evaluate the performance of a trained decoder (as shown in Figure 6 and Table A.2), we fine-tune a gemma-3-12b-it decoder models based on a frozen gpt-5 encoder on both EEDI and XES3G5M.
+Data Construction
+We train on 800 (resp.
+400) students from the training split of the XES3G5M Filtered (resp.
+EEDI Filtered) dataset, which amounts to about 80% of the data, leaving 200 test students.
+For each student interaction history $H_{t}$, we first generate a bottleneck summary $\tilde{\mathcal{S}}=f_{\theta}(H_{t})$ using the frozen gpt-5 encoder.
+We then pair this summary with the ground-truth binary correctness labels for the $|\mathcal{Y}|=4$ target questions associated with that student, yielding a training dataset of $N=3,200$ (resp.
+$N=1,600$) training examples.
+Training Procedure
+For each bottleneck $\tilde{\mathcal{S}}$ and question $q$, the model is prompted to predict the correctness of the student via the decoder prompt shown in Listing LABEL:lst:base_prompt_dec.
+We then train with SFT on the ground truth answers using LoRA (Hu et al., 2022) on a single NVIDIA RTX 6000 48GB GPU.
+We use a train/validation split of 90/10 and train with early stopping on validation accuracy computed after each epoch.
+The training configuration:
+  • Model: google/gemma-3-12b-it
+  • LoRA Configuration: Rank $r=16$, Alpha $\alpha=16$. Target modules: [q_proj, k_proj, v_proj, o_proj].
+  • Optimization: AdamW optimizer with a learning rate of $1\times 10^{-4}$.
+  • Batch Size: Effective batch size of 8 (per-device batch size of 2 with 4 gradient accumulation steps).
+  • Duration: The model was trained for up to 5 epochs with early stopping after the first decreasing validation accuracy.
+B.3 KT/CD Models
+Cognitive Diagnosis models
+We evaluate the following 5 Cognitive Diagnosis models:
+  • IRT (Rasch, 1993)
+  • MIRT (Reckase, 2006)
+  • DINA (Junker and Sijtsma, 2001)
+  • KaNCD (Wang et al., 2022)
+  • NeuralCDM (Wang et al., 2022)
+We use the implementation from the EduCDM library (bigdata-ustc, 2021). To make sure there is enough data per student to train on, we filter out students with <10<10 interactions in each dataset.
+CD baselines CD models were implemented using the EduCDM library (bigdata-ustc, 2021).
+Unlike KT and LBMs, CD models learn static, student-specific parameters (e.g., proficiency embeddings) which requires every student to be present during training.
+Consequently, they cannot be evaluated on previously unseen students in a zero-shot manner.
+To ensure a fair comparison, we evaluate CD models using a held-out set of $|\mathcal{Y}|=4$ interactions for each student.
+In this setup, the model “reads” the student’s history by optimizing student-specific parameters via gradient descent.
+We consider this functionally equivalent to how KT and LBMs process the same history “in-context” (via RNN states or attention) to adapt to a new student.
+Therefore, despite the protocol difference, all models use the same historical information $\mathcal{X}$ to predict $\mathcal{Y}$.
 Knowledge Tracing models
-We evaluate the following four Knolwedge Tracing models:
-  • DKT [20]
-  • DKVMN [34]
-  • AKT [6]
-  • SAKT [19]
-For each one of these models, we use the pyKT implementation [15] with default hyperparameters.
-D.2 Datasets
-D.2.1 Synthetic
+We evaluate the following 9 Knowledge Tracing models:
+  • DKT (Piech et al., 2015)
+  • DKVMN (Zhang et al., 2017)
+  • SAKT (Pandey and Karypis, 2019)
+  • GKT (Nakagawa et al., 2019)
+  • Deep IRT (Yeung, 2019)
+  • AKT (Ghosh et al., 2020)
+  • SAINT (Choi et al., 2020)
+  • SimpleKT (Liu et al., 2023a)
+  • QIKT (Chen et al., 2023)
+KT baselines KT models are evaluated using the pyKT library (Liu et al., 2022), modified to compute accuracy on the same $|\mathcal{Y}|=4$ target window used for LBMs.
+We use the default hyperparameters from pyKT for all models.
+Test accuracy is computed on a set of $200$ unseen test students.
+B.4 Datasets
+B.4.1 Synthetic
 Our synthetic dataset simulates students answering basic arithmetic problems.
@@
   • fails any operation involving the number $x$ with $x\sim\mathcal{U}(6,9)$;
-  • fails whenever an operand $>10$;
+  • fails whenever an operand >10>10;
   • always rounds division results down;
@@
   • Questions answered per student: $210$.
-Figure D1 shows the histogram of the number of misconceptions per student across the $2000$ students of the dataset.
-Figure D1: Distribution of the number of misconceptions per student in the Synthetic dataset.
-D.2.2 Eedi
-The Eedi dataset is analogous to the one publicly shared via the NeurIPS 2020 Education Challenge [30] organised by Eedi, but additionally includes the text of each question.
-Although many trajectories in the Eedi dataset span several days or months, we retain only single-session sequences to satisfy the quasi-static-knowledge assumption:
-  • minimum response time: $5\,\mathrm{s}$;
+Figure B1 shows the histogram of the number of misconceptions per student across the $2000$ students of the dataset.
+Figure B1: Distribution of the number of misconceptions per student in the Synthetic dataset.
+B.4.2 Eedi
+The Eedi dataset is analogous to the one publicly shared via the NeurIPS 2020 Education Challenge (Wang et al., 2020) organised by Eedi, but additionally includes the text of each question.
+While the exact version of the dataset used in this work is not available publicly, a very similar version including question texts is available via the "Eedi - Mining Misconceptions in Mathematics" Kaggle Competition (King et al., 2024).
+Preprocessing
+Although many trajectories in the Eedi dataset span several days or months, we retain only single-session sequences to satisfy the constant knowledge state assumption - a common assumption in Cognitive Diagnosis (Wang et al., 2024).
+We compute sessions of interactions by grouping answers that are within the follow criteria:
+  • minimum response time: $5\,\mathrm{s}$, to avoid random answers;
   • maximum gap between answers: $3\,\mathrm{min}$;
   • minimum trajectory length: $40$ questions.
-This filtering yields a total of $623$ individual trajectories with 40+ questions.
-D.3 Results
-D.3.1 Table 1
-Synthetic dataset
-  • $|\mathcal{X}|=50$
-  • $|\mathcal{Y}|=4$
-  • Test students: $200$
-  • Bottleneck size: $128$ tokens
-  • Chain-of-thought prompting: No
-Eedi dataset
-  • $|\mathcal{X}|=30$
-  • $|\mathcal{Y}|=4$
-  • Test students: $100$
-  • Bottleneck size: $512$ tokens
-  • Chain-of-thought prompting: No
-D.3.2 Figure 4
-Same set-up as Table 1; only the bottleneck size varies (see $x$-axis).
-D.3.3 Figure 5
+This filtering yields the "EEDI Filtered" dataset with a total of $623$ individual trajectories with 40+ questions.
+B.4.3 XES3G5M
+The XES3G5M dataset contains student interaction data from a large-scale online mathematics learning platform (Liu et al., 2023b).
+The original dataset has question and construct text in Chinese.
+We use English translation from Ozyurt et al.
+(2024) to also run on an English version of the dataset.
+Preprocessing
+The XES3G5M preprocessing follows a similar approach to Eedi, with adjusted parameters to accommodate the characteristics of this dataset.
+We compute sessions of interactions by grouping answers that meet the following criteria:
+  • minimum response time: $5\mathrm{s}$, to avoid random answers;
+  • maximum gap between answers: $10\mathrm{min}$;
+  • minimum trajectory length: $34$ questions.
+The increased maximum gap between answers (10 minutes vs.
+3 minutes for Eedi) is to ensure that a sufficient number of sessions are available for training.
+More stringent filtering make it easier to satisfy the constant knowledge state assumption, but might not produce enough data points for effective training of KT and CD models.
+This filtering yields the "XES3G5M Filtered" dataset with a total of $1,245$ student-sessions trajectories with 34+ questions, across $996$ individual students.
+B.5 Experiment Parameters
+B.5.1 Table A.2, Figure 6A, Table 2, Figure 7
+Table B1 summarize experimental settings used for Table A.2, Figure 6, Table 2 and Figure 7.
+Runs are aggregated across N=3 for LBMs/Direct LLMs and N=10 for KT/CD models.
+For Figure 7 the bottleneck size varies along the plot’s x axes.
+Table B1: Experimental settings for the different datasets.
+Synthetic Eedi (Filtered) XES3G5M (Filtered)
+$|\mathcal{X}|$ 50 30 30
+$|\mathcal{Y}|$ 4 4 4
+# Train students 400 800
+# Test students 200 100 200
+Bottleneck size 128 tokens 512 tokens 512 tokens
+CoT prompting No No No
+Statistical Test for Table 2
+Table 2 shows in bold the models and datasets for which the LBM variant has an accuracy of no more than 2% below the Direct baseline.
+We use Welch’s t-test with the null hypothesis: $\texttt{accuracy}_{LBM}>=\texttt{accuracy}_{direct}-0.02$‘, and report in bold runs where $p<0.05$.
+B.5.2 Figure 6B
 Latent Bottleneck Models (LBMs)
-  • $|\mathcal{X}|=50$, $|\mathcal{Y}|=4$
+  • $|\mathcal{X}|=N$, $N\in\{5,10,20,50\}$, $|\mathcal{Y}|=4$
   • Test students: $200$
   • Bottleneck size: $256$ tokens
-  • Chain-of-thought prompting: Yes
-  • Encoder & decoder: identical (varies by row in the table)
-Knowledge-tracing baselines DKT [20], DKVMN [34], AKT [6], and SAKT [19] are evaluated using the pyKT implementation [15], modified to compute accuracy on $|\mathcal{Y}|=4$ questions, in order to be comparable with our LBM models.
-For each method we keep the pyKT default hyperparameters.
-D.3.4 Figure 6
+  • Chain-of-thought prompting: Yes; reasoning_effort=minimal for GPT-5
+  • Encoder & decoder: same backbone LLM (varies by row in the table)
+B.5.3 Training the encoder (Figure 5)
   • Dataset: Synthetic
@@
 Training hyper-parameters.
-[…]
+  • $\text{Batch size}:5$
   • $G=5$
-  • Learning $\text{Learning rate}=1\times 10^{-4}$
+  • $\text{Learning rate}=1\times 10^{-4}$
   • $\beta=0.04$
-  • Optimiser: Adam (adamw_torch, default settings)
+  • Optimizer: Adam (adamw_torch, default settings)
   • LoRA configuration
@@
     • Dropout: $0.05$
-D.3.5 Encoder–Decoder Variants
-Identical to the Table 1 set-up, except for the choice of encoder and decoder LLMs.
-D.3.6 Difficulty Variations
-Synthetic students are stratified by the number of misconceptions held: 00, $1$, $2$, or $3+$.
-D.3.7 Steering Experiments
+B.5.4 Encoder–Decoder Variants
+Identical to the Table A.2 set-up, except for the choice of encoder and decoder LLMs.
+B.5.5 Steering Experiments
 Same dataset and model parameters as for Figure 5.
 Training hyper-parameters.
-[…]
+  • $\text{Batch size}:4$
   • $G=4$
-  • Learning $\text{Learning rate}=1\times 10^{-4}$
+  • $\text{Learning rate}=1\times 10^{-4}$
   • $\beta=0.04$
-  • Optimiser: Adam (adamw_torch, default settings)
+  • Optimizer: Adam (adamw_torch, default settings)
   • LoRA configuration
@@
     • Dropout: $0.05$
-D.4 Prompts
+B.6 Prompts
 Below are the different prompts used to query LLMs in our experiments.
@@
 11 Keep your summary under {max_words} words.
+12
+13# with CoT
+14cot_instruction: |
+15 Think step by step and lay out your reasoning before you write the final summary. Then, enclose the final summary in <info>...</info>.
+16# without CoT
+17cot_instruction: |
+18 Enclose your entire summary in <info>...</info> and do not include anything else.
+19
+20input_text: |
+21 The student answered the following questions:
+22 {question_1_text}
+23 ...
+24 {question_n_text}
+25
+26# with construct information
+27question_i_text: |
+28 Question {question_txt, question_ID}, with construct{construct_txt, construct_ID}: answered {correctness}.
+29
+30# without construct information
+31question_i_text: |
+32 Question {question_txt, question_ID}: answered {correctness}.
 Listing 1: Base Prompt for the Encoder.
@@
 4
-5 {y_prompt}
+5 Predict whether the student will answer {new_question} correctly or not.
+6 Answer with "Yes" or "No" and nothing else.
 Listing 2: Base Prompt for the Decoder.
@@
 13 Keep your summary under {max_words} words.
-Listing 3: Base Prompt for the Encoder when steering for mentioning "misconceptions" (Section B.4.1).
-D.5 Code and hardware
+Listing 3: Base Prompt for the Encoder when steering for mentioning "misconceptions" (Section A.4.2).
+B.7 Code and hardware
 Experiments were ran on four NVIDIA A100 GPUs with 80GB VRAM each, and 880GB of total RAM.
+Appendix C Systematic Evaluation via LLM-as-a-judge
+C.1 Method
+We use LLM-as-a-judge (Li et al., 2025) with GPT-5 to systematically evaluate generated summaries on the Synthetic dataset.
+Specifically, for a given student in the Synthetic dataset we provide GPT-5 with the ground truth knowledge state and a summary to be evaluated.
+The judge LLM is blinded to the model that produced the summary.
+We then prompt the model to assess the summary across the following dimensions:
+  • Misconception detection: for each misconception in the student knowledge state, assess how whether the summary captures it.
+Score: 0,1,2, 0 for not mentioned, 1 for partial mention (implied or vaguely hinted at), 2 for explicitly stated.
+  • Misconception false positives: report any instance of misconception stated in the summary that do not appear in the ground truth knowledge state.
+  • Construct-level accuracy: for each construct, determine whether the construct’s true mastery was accurately captured by the summary. Score: 0,1, 0 for incorrect or missing, 1 for correct.
+  • Overall score: generally assess the alignment of the summary with the ground truth. Score: 1 to 5, from major mismatch (1) to very strong alignment (5).
+  • Confidence calibration: evaluate whether the summary expresses uncertainty appropriately. Score: 1 to 3, 1 for under-confident and vague, 2 for well-calibrated, 3 for over-confident.
+  • Specificity: evaluate how precise the summary is. Score: 1 to 3, 1 for very generic and vague, 2 for moderately specific, 3 for highly specific.
+C.2 Evaluation examples
+Table LABEL:app:llm_eval_examples shows the result of the evaluation for different summaries generated for student #1585.
+Table C1: Comparison of generated student summaries and GPT-5-based LLM-as-a-judge evaluation (Synthetic dataset, student #1585).
+Keys: Q.: Overall Score (1-5, higher is better).
+Constructs (0/1): Add (Addition), Sub (Subtraction), Mul (Multiplication), Div (Division).
+Misconceptions (0-2, 0=missed, 1=implied, 2=explicitly noted): Div R (Rounds down division), Add C (Forgets carry), Neg (Fails negatives), x7 (Fails x7).
+FP: False Positives count.
+Cal: Confidence Calibration (1-3, 1=under-confident, 2=appropriately calibrated, 3=over-confident).
+Spec: Specificity (1-3, higher is more specific).
+Model Generated Summary Q. Constructs (0/1) Misconceptions (0-2) FP Cal Spec
+(1-5) Add Sub Mul Div Div R Add C Neg x7
+Ground Truth The student has mastered all skills (Add, Sub, Mul, Div) but has systematic misconceptions: Always rounds down division; Forgets to carry in addition; Fails with negative numbers; Fails multiplication involving 7.
+- 1 1 1 1 2 2 2 2 - - -
+gpt-5 Student is strong in basic addition (mostly correct for sums up to ~25, but missed 3+8, 8+12, 6+15, 15+8).
+Subtraction: competent with small positive results (e.g., 10-4, 14-12, 15-13, 5-1), but frequently incorrect on negatives/zero-minus cases (3-6, 0-12, 3-15, 3-9, 0-15, 5-11, 4-9, 3-7, 1-15).
+Multiplication: generally solid (facts like 2×2, 4×1, 4×5, 6×5, 6×10, 9×8 correct) but missed 5×7.
+Division (rounded): good with whole-number quotients (÷1, 10÷2, 4÷2, 6÷1) and 6÷5, but struggles with non-integer cases (8÷5, 6÷4, 5÷3).
+5 1 1 1 1 2 1 2 1 0 2 3
+gpt-4o The student consistently answers addition (construct ID 0) and multiplication (construct ID 2) questions correctly, reflecting strong competence in these areas.
+They show varying performance in subtraction (construct ID 1), struggling with negative results.
+In division (construct ID 3), they often answer correctly when the result is a whole number but struggle when rounding is required.
+Their overall performance indicates proficiency in basic arithmetic, particularly addition and multiplication, with some difficulty in questions involving subtraction with negative results and non-whole arithmetic operations.
+4 1 1 1 1 1 0 2 0 1 2 2
+gemma-3-12b The student demonstrates strong addition and multiplication skills (constructs 0 & 2) with occasional errors.
+Subtraction (construct 1) is a significant weakness, showing frequent mistakes across various numbers.
+Division with rounding (construct 3) is inconsistent; the student excels with simple divisions but struggles with others.
+Performance is notably worse with negative results or larger numbers in subtraction and occasionally addition.
+The student appears to grasp basic arithmetic operations but lacks fluency and accuracy in subtraction.
+2 0 0 0 0 1 1 2 0 2 3 2
+Qwen2.5-3B The student has strong accuracy in arithmetic operations like addition and multiplication, particularly when constructs are 0, 2, or 3.
+However, they struggle with subtraction and division, especially when rounding is involved (constructs 1 and 3).
+2 0 0 1 0 1 0 0 0 1 3 2
+gpt-4o-mini The student struggles with basic subtraction problems, having answered incorrectly on many related questions.
+They demonstrate understanding of addition and multiplication, scoring correctly on multiple questions in those areas.
+The student has shown some success in division, rounding, but struggles with specific datasets.
+They can solve problems involving simple addition, like 10 - 4, but make errors with negative results and more complex subtraction like 3 - 15.
+Overall, their strengths lie in addition and multiplication, while they need improvement in subtraction and certain division problems.
+2 0 0 1 0 1 0 2 0 2 3 2
+Qwen2.5-7B The student struggles with subtraction (construct ID 1) and division (construct ID 3), while performing well in addition (construct ID 0) and multiplication (construct ID 2).
+Correct answers include 10/2 (5), 11+10 (21), 0+6 (6), 6/5 (1), 12+1 (13), 4*5 (20), 9*4 (36), 15-13 (2), 13+12 (25), 6/1 (6), 8*1 (8), 5*7 (35), 2+13 (15), 9*8 (72), 9*8 (72), 4/2 (2).
+1 0 0 0 0 0 0 0 0 0 3 2
+gemma-3-27b The student reliably performs addition and multiplication with smaller positive integers (constructs 0 & 2).
+They struggle with subtraction and division, especially when results are negative or not whole numbers (constructs 1 & 3).
+Accuracy decreases with larger numbers in subtraction and addition.
+Division rounding is often incorrect.
+Performance on construct 0 is ~85% correct, construct 1 is ~40% correct, construct 2 is ~75% correct, and construct 3 is ~50% correct.
+1 0 0 0 0 1 0 2 0 2 3 2
+C.3 Prompts
+We use the following system prompt and prompt template to instruct and query the judge model.
+1system_prompt: |
+2 You are an expert evaluator assessing whether a model-generated summary ("bottleneck") faithfully reflects a synthetic students knowledge state.
+3
+4 The synthetic students follow this rule:
+5 - A mastered construct is always answered correctly *unless* a specific misconception applies.
+6 - A non-mastered construct is always answered incorrectly.
+7 Therefore, misconceptions tied to non-mastered constructs effectively never appear in the data and should not be expected in the bottleneck.
+8
+9 You will evaluate the bottleneck on **six components**:
+10 1. Misconception capture (0-2 each)
+11 2. Misconception false positives (list of str)
+12 3. Construct-level accuracy (0-1 each)
+13 4. Overall quality (1-5)
+14 5. Confidence calibration (1-3)
+15 6. Specificity (1-3)
+16
+17 # INPUT
+18 ...
+19 # Bottleneck
+20 {bottleneck_text}
+21
+22 # Ground Truth Constructs
+23 addition: mastered/not mastered
+24 subtraction: mastered/not mastered
+25 multiplication: mastered/not mastered
+26 division: mastered/not mastered
+27
+28 # Ground Truth Misconceptions
+29 Misc 0: {text}
+30 Misc 1: {text}
+31 ...
+32
+33 # EVALUATION RULES
+34
+35 1. *Misconception capture (0-2)*
+36 For each misconception ("Misc 0", "Misc 1", etc.), judge how well the bottleneck captures it.
+37 Use the following scale:
+38 - 0 = Not captured -- the bottleneck does not mention or imply this misconception.
+39 - 1 = Partially captured -- indirectly implied or vaguely hinted at.
+40 - 2 = Clearly captured -- explicitly stated or strongly described.
+41 Output:
+42 "misc_scores": {"0": <0/1/2>, "1": <0/1/2>, ...}
+43
+44 2. *Misconception false positives*
+45 A false positive occurs when the bottleneck **mentions or implies a systematic misconception that is NOT present in the ground truth list**.
+46
+47 List each instance of such incorrect misconception by extracting a short excerpt of the relevant text from the bottleneck. If none are found, return an empty list.
+48 Output:
+49 "misc_false_positive": ["...", "..."]
+50
+51 3. *Construct-level accuracy (0-1)*
+52 There are four constructs: addition, subtraction, multiplication, division.
+53 Score whether the bottleneck correctly reflects the students mastery status, **taking misconceptions into account**.
+E.g., if a construct is mastered except when a misconception applies, statements like "good at X except in specific cases" count as correct.
+54 Score 1 if the construct is accurately described; score 0 if inaccurate or missing.
+55 Output:
+56 "construct_scores": {"addition": <0/1>, "subtraction": <0/1>, "multiplication": <0/1>, "division": <0/1>}
+57
+58 4. *Overall quality (1-5)*
+59 Holistic score of how well the bottleneck aligns with the entire ground truth.
+60 - 1 = Major mismatch
+61 - 2 = Several substantial errors
+62 - 3 = Mixed; partial alignment
+63 - 4 = Mostly correct with minor issues, such as rare misconceptions
+64 - 5 = Very strong alignment
+65 Output:
+66 "overall_score": <1-5>
+67
+68 5. *Confidence calibration (1-3)*
+69 Evaluate whether the bottleneck expresses uncertainty appropriately:
+70 - 1 = Underconfident (hedges excessively; too vague or noncommittal)
+71 - 2 = Well-calibrated (claims are proportional, avoids strong unwarranted assertions)
+72 - 3 = Overconfident (makes strong claims not justified by the data)
+73 Output:
+74 "confidence_calibration": <1-3>
+75
+76 6. *Specificity (1-3)*
+77 Evaluate how specific or generic the bottleneck is:
+78 - 1 = Very generic; vague statements without detail
+79 - 2 = Moderately specific; some detail, but missing precision
+80 - 3 = Highly specific; clearly articulated patterns that closely match the ground truth
+81 Output:
+82 "specificity": <1-3>
+83
+84 # OUTPUT FORMAT
+85
+86 Return only a dictionary in the following structure:
+87 {
+88 "misc_scores": {"0": 2, ...},
+89 "misc_false_positive": ["’struggles with rounding’, ...]
+90 "construct_scores": {"addition": 1, "subtraction": 0, "multiplication": 1, "division": 1},
+91 "overall_score": 4,
+92 "confidence_calibration": 2,
+93 "specificity": 3,
+94 }
+95
+96 Do not include anything else in your answer.
+97
+98user_prompt_template: |
+99 # Bottleneck
+100 {bottleneck}
+101
+102 # Ground Truth Constructs
+103 {constructs}
+104
+105 # Ground Truth Misconceptions
+106 {misconceptions}
+Listing 4: Prompts used in LLM-as-a-judge evaluation.
+Appendix D Steerability of the Estimated Knowledge State
+A key advantage of LBMs is the ability for humans to interact with the model to steer the estimated student knowledge states and complement the model with additional information.
+Here, we further discuss the three mechanisms for human-model interaction in the LBM framework mentioned in Section 3.4.
+Prompt engineering the encoder.
+The most straightforward approach is directly shaping how the encoder generates summaries through prompt engineering, for example via system instructions or in-context examples (Brown et al., 2020).
+By modifying the instruction prompt given to the encoder, educators can influence the format, emphasis, and level of detail in the generated knowledge state summaries.
+For instance, instructing the encoder to highlight specific types of misconceptions or to focus on particular subject areas can yield more targeted summaries.
+Examples of good and bad knowledge states can also be provided to the encoder to steer its behavior via in-context learning.
+Steering via reward signals during training.
+When training the encoder, human preferences can be incorporated through the reward function by using the $\Omega(S)$ term from Eq.
+1.
+This allows for systematic enforcement of desirable summary properties across the model’s outputs.
+For example, if educators find that explicit enumeration of misconceptions is particularly valuable for intervention planning, the reward function can be designed to favor summaries that consistently identify and articulate student misconceptions – as demonstrated Figure A.4.2.
+Augmenting with student-specific information.
+Perhaps the most powerful form of interaction involves supplementing the model with additional student-specific information not present in the observed interaction data $X$.
+This can occur in two ways:
+  • Augmenting encoder input: Educators can provide supplementary information alongside the observed interactions $X$, such as notes about recent classroom activities not yet reflected in assessment data, or observations about a student’s learning process not captured in their answers.
+  • Modifying the generated summary: After the encoder produces a knowledge state summary $S$, educators can directly edit this summary based on their domain expertise and student-specific knowledge before passing it to the decoder.
+For example, a teacher who noticed a student struggling with negative numbers during an in-class exercise could add this observation to the generated summary, even if the available assessment data contains few questions involving negative numbers.
+Appendix E Computation Cost Analysis
+Table LABEL:tab:comp_runtime compares the average wall-clock runtime of different model types across our experiments.
+LBMs take up to 10x more time to run than traditional methods due to LLM inference costs, and roughly 2x longer than the Direct variant (as it requires inference on both the encoder and decoder).
+Despite higher computational costs, LBMs offer unique qualitative interpretability, nuanced insights and zero-shot capabilities that justify such overhead for educational applications where interpretability is key and training data is limited.
+Table E1: Average computation time across (in seconds) of each model across our experiments, aggregated by model type (mean±std).
+KT models are run on an NVIDIA RTX 6000 48GB, the CD, Direct and LBM models are run on an NVIDIA A100 80GB.
+Synthetic EEDI (Filtered) XES3G5M (Filtered)
+Model Type
+CDM (Training + Inference) 24.8 ± 22.5 69.3 ± 70.5 131.0 ± 195.9
+KT (Training + Inference) 180.5 ± 333.8 149.0 ± 237.1 242.2 ± 498.9
+Direct (Inference) 268.6 ± 262.4 629.0 ± 863.0 479.3 ± 371.8
+LBM (Inference) 445.0 ± 629.2 1384.3 ± 2091.1 1352.7 ± 2091.7
+Appendix F Extended Related Work
+This section provides an extended review of the literature related to our proposed approach, spanning traditional knowledge tracing, recent advances in LLM-based student modeling, and concept bottleneck models (CBMs), and text summarization models.
+F.1 Cognitive Diagnosis
+Cognitive Diagnosis Models (CDMs) aim to infer a student’s latent knowledge state from their observed responses to test questions (Wang et al., 2024).
+F.1.1 Psychometrics-based CD models
+Classical models for Cognitive Diagnosis originate from psychometrics, including Item Response Theory (IRT) and its multidimensional variant (MIRT), which model continuous scores of knowledge proficiency using logistic functions (Rasch, 1993; Reckase, 2006).
+The DINA model estimates binary mastery variables of knowledge concepts, assuming that students must master all required skills to answer correctly, while accounting for slips and guesses (Junker and Sijtsma, 2001).
+Other variants have been proposed with different assumptions, such as DINO (Templin and Henson, 2006) which considers that students will correctly answer the item if at least one required knowledge concept is mastered.
+A critical input for these methods is the Q-matrix, which describe which knowledge concept is required for each question.
+F.1.2 Deep Learning-based CD models
+More recent deep learning approaches offer greater flexibility in modeling complex relationships.
+The Neural Cognitive Diagnosis Model (NCDM) uses neural networks to learn interaction functions between student proficiency vectors and item characteristics (Wang et al., 2022).
+Extensions include Kernel-based Neural Cognitive Diagnosis (KaNCD), which models latent associations between knowledge concepts (Wang et al., 2022), and Knowledge-Sensitive Cognitive Diagnosis (KSCD), which learns intrinsic relations between knowledge concepts from student responses (Ma et al., 2022).
+Graph neural networks have also been incorporated, with frameworks like RCD capturing relationships between students, questions, and knowledge concepts (Gao et al., 2021).
+Recent encoder-decoder architectures, such as ID-CDF, enable inductive diagnosis by directly encoding student responses into ability vectors (Li et al., 2024b).
+While these deep learning models provide enhanced predictive power and can handle diverse data types, they still typically operate within predefined knowledge concept frameworks and are limited to quantitative estimates of skill mastery.
+F.2 Knowledge Tracing
+Compared to Cognitive Diagnosis which assumes a constant knowledge state, Knowledge Tracing methods aim at estimating evolving knowledge states as students answer questions.
+We similarly review Knowledge Tracing methods (Shen et al., 2024), as well as recent efforts to enhance the interpretability of KT models.
+F.2.1 Deep Knowledge Tracing
+Deep learning-based approaches like Deep Knowledge Tracing (DKT) (Piech et al., 2015), Dynamic Key-Value Memory Networks (DKVMN) (Zhang et al., 2017) and Attentive Knowledge Tracing (AKT) (Ghosh et al., 2020) uses neural networks or attention-based architecture to learn contextual representation of questions and student knowledge states.
+Despite their strong predictive performance, these models represent a student’s knowledge state as an abstract, high-dimensional latent vector, which poses significant challenges in interpretability and actionable feedback for educators.
+F.2.2 Interpretable Knowledge Tracing
+Several recent works have proposed more interpretable Knowledge Tracing methods.
+Bayesian methods for Knowledge Tracing
+Early KT approaches used structured models where learned parameters allow for direct interpretations.
+For example, in Bayesian KT (Corbett and Anderson, 1994) and dynamic Bayesian KT (Käser et al., 2017), the latent variables learned represent the student’s evolving proficiency across knowledge concepts (KCs).
+Interpretable Knowledge Tracing (IKT) (Minn et al., 2022) introduces a causal probabilistic student model based on skill mastery, ability profiles, and problem difficulty.
+While this approach provides clearer connections between model components and predictions, interpretability remains tied to quantitative skill mastery estimates.
+PSI-KT (Zhou et al., 2024) similarly learns a student’s latent knowledge state over KCs via a probabilistic hierarchical state-space model and psychology-inspired learning dynamics.
+While these methods provide interpretable latent states, they remain limited to estimating the student’s proficiency over KCs.
+IRT-based Knowledge Tracing
+Other works combine deep learning architectures with classical item-response theory to make latent representation interpretable.
+Deep-IRT (Yeung, 2019) leverages the powerful abilities of deep learning architectures by training a DKVMN model to estimate student ability on each KC and item difficulty and predict future answers via an IRT model.
+QIKT (Chen et al., 2023) employs IRT functions as the final prediction layer, combining question-centric knowledge acquisition, knowledge mastery scores, and knowledge application scores through encoder modules.
+Despite meaningful latent representations, the diagnostic output of these methods remains constrained to proficiency estimates over KCs.
+Learned Questions Relationships
+HGKT (Tong et al., 2022) addresses limitations of concept-based proficiency by modeling hierarchical relationships between questions and introducing problem schemas as additional links between questions.
+Problem schemas are discovered through hierarchical clustering of question embeddings via BERT encodings, providing a means of grouping questions orthogonal to knowledge concepts and therefore enabling more detailed diagnostic reports than at the concept proficiency level.
+However, the interpretability of these learned schemas requires post-hoc interpretation using TextRank to infer schema descriptions from clusters of questions.
+While this approach moves beyond simple concept proficiency, it still relies on learned embeddings that cannot directly articulate student reasoning patterns without post-hoc manipulations.
+Explainable subsequences
+Explainable subsequences provide an alternative to interpretability via concept proficiency by identifying which past questions are most relevant for predicting future responses.
+For example, Li et al.
+(2023) proposes a genetic algorithm to identify explainable subsequences in student interaction histories better than with standard Deep Learning explanation methods such as Shapley values or gradient-based saliency maps.
+While this allows for a different kind of interpretability from concept mastery, the explanations remain at the level of question relevance rather than underlying reasoning processes.
+This approach can indicate which questions matter but cannot explain why they matter or what misconceptions they reveal.
+Option Tracing
+Option Tracing (Ghosh et al., 2021) moves beyond binary correctness by modeling which specific option a student selects in multiple-choice questions, enabling finer-grained analysis of misconceptions through patterns in distractor choices.
+Similarly, Park et al.
+(2024) leverages MCQ responses and concept maps to disentangle student understanding at the concept level.
+While their approach is motivated by misconception detection, their IRT-based predictions remain grounded in concept-level proficiency prediction and do not explicitly validate misconception identification.
+F.2.3 LLM-Based Knowledge Tracing
+Recent studies have begun integrating Large Language Models (LLMs) into the KT framework.
+For example, (Li et al., 2024a) demonstrated that LLMs are able to make sensible predictions about student responses when prompted with adequate information.
+Other works (Lee et al., 2024; Kim et al., 2024) have studied how LLMs can help mitigate the cold-start problem compared to traditional KT approaches, while (Wang et al., 2025) demonstrated state-of-the-art performances in KT by combining LLMs with sequence interaction models.
+However, these methods generally remain opaque: they either treat the LLM as a black-box, or rely on model-generated explanations that are susceptible to hallucination (Bender et al., 2021).
+KCQRL (Ozyurt et al., 2024) leverages language models to improve the embedding of any deep learning KT method by encoding semantic information about question content, while KCD (Dong et al., 2025) refines a cognitive diagnosis model using the general knowledge an LLM.
+While these methods can improved the predictive accuracy of KT and CD models, their interpretability remains limited to knowledge concept proficiency.
+F.3 Comparison of LBMs to KT and CD
+Table F1 gives a high-level comparison of LBMs and KT/CD models.
+We acknowledge that this summary table is a simplification of the KT and CD fields, as each contains many works that have been proposed to tackle these individual limitations (for example, using textual question contents to improve embeddings (Ozyurt et al., 2024), or using specialized Multiple Choice Questions to reveal misconceptions via KT-type methods(King et al., 2024)).
+Nevertheless, it clarifies our paper’s similarities with these two fields, while contrasting key differences that enable LBMs to address educational scenarios where neither traditional KT nor CD methods are well-suited —particularly for misconception identification or when working with limited data.
+Table F1: Comparison of Cognitive Diagnosis, Knowledge Tracing and our proposed Language Bottleneck Models.
+Aspect Knowledge Tracing (KT) Cognitive Diagnosis (CD) Language Bottleneck Models (LBMs) Interpretability Capabilities Interpretable Output Quantitative proficiency across knowledge concepts Quantitative proficiency across knowledge concepts Qualitative text summaries Misconception Detection No No Yes Requires Predefined or Inferred Concepts for Interpretability Yes Yes No Data Requirements Primary Input Modality Question and/or Construct IDs Question and Construct IDs (Q-matrix) Any textual information Training Data Requirements High High Low/Zero-shot Modeling Characteristics Knowledge State Assumption Dynamic Static Static Human-in-the-Loop Limited Limited High (steerable & editable) Fundamental Distinctions Core Question Addressed “Can we predict a student’s responses over time?” “Can we quantitatively estimate a student’s proficiency across concepts?” “Can we qualitatively estimate a student’s knowledge state, including knowledge concepts and misconceptions?” Case Study (Figure 3) Similar to CD Proficiency vector: KC1 (Add): 0.59, KC2 (Sub): 0.53, KC3 (Mul): 0.76, KC4 (Div): 0.23 “…excels at addition and multiplication… Subtraction is a weakness… Multiplication by 6 or 7 seems to be a specific area of difficulty…”
+F.4 Concept Bottleneck Models
+Concept Bottleneck Models (CBMs) (Koh et al., 2020) improve interpretability through human-understandable concept activations as intermediates, with extensions exploring unsupervised concept learning (Oikarinen et al., 2023), test-time interventions (Shin et al., 2023), and theoretical analyses of concept set design (Luyten and van der Schaar, 2024).
+However, CBMs typically rely on finite predefined concept sets, limiting their applicability to complex tasks like knowledge tracing.
+Recently, (Yamaguchi and Nishida, 2024) introduced Explanation Bottleneck Models (XBMs), which use textual rationales as intermediates for vision classification, justifying a single known label per input.
+While our Language Bottleneck Models (LBMs) adopt this language bottleneck concept, they differ fundamentally.
+Unlike XBMs’ instance-specific, task-specific rationales, LBM summaries aim to capture an implicit knowledge states—as emphasized by our inverse problem formulation—and generalize to future, unknown questions.
+These requirements necessitate holistic, adaptable summaries rather than XBMs’ task-specific rationales, leading us to introduce Language Bottleneck Models (LBMs) as a distinct framework with broader applicability.
+F.5 Text Summarization Models
+Recent text summarization models such as BART (Lewis et al., 2019), T5 (Raffel et al., 2020), and PEGASUS (Zhang et al., 2020) effectively produce concise summaries based on explicitly available textual content.
+However, these approaches differ fundamentally from knowledge tracing, where summaries must infer implicit student knowledge states not directly observable in the input.
+Standard summarization metrics (e.g., ROUGE, BLEU) rely on explicit reference summaries, making them unsuitable for evaluating latent knowledge inference tasks.
+In contrast, our Language Bottleneck Models generate textual summaries of the student’s implicit knowledge state, optimized for predictive accuracy on downstream questions rather than syntactic overlap with the observed input.
+Appendix G Extended Discussion
+G.1 Extended Discussion
+Parallel with Kolmogorov Complexity
+The inverse problem formulation of Knowledge State Modeling draws a natural parallel with algorithmic information theory.
+The Kolmogorov complexity $K(x)$ of a string $x$ is defined as the length of the shortest program that outputs $x$ when run on a universal Turing machine.
+Traditional KT methods focus on learning statistical regularities between questions and responses, without looking for parsimonious explanations for observed behavior.
+In contrast, LBMs explicitly search for minimal natural language descriptions that can both reconstruct past interactions and predict future responses.
+The knowledge state summary $\mathcal{S}$ is akin to the minimal program, and the decoder LLM functions is akin to the universal interpreter that unpacks $\mathcal{S}$ to generate the observed question-answer patterns.
+This connection suggests that effective knowledge state summaries should capture the algorithmic essence of student behavior—the underlying "program" of knowledge and misconceptions that generates observable responses—rather than merely fitting surface-level patterns.
+While true Kolmogorov complexity is uncomputable, LBMs approximate this ideal through the natural language bottleneck constraint, encouraging summaries that balance compression with predictive power.
+Computational Architecture and Design Choices
+The encoder-decoder architecture reflects a deliberate separation of concerns: extraction versus interpretation of knowledge states.
+Our experiments demonstrate that these two tasks have different intrinsic complexity, with encoding (summary generation) being more challenging than decoding (prediction from summaries).
+This finding has practical implications for model selection and computational resource allocation.
+The use of GRPO for encoder training represents a novel application of reinforcement learning to interpretable AI, where the reward signal directly measures the downstream utility of explanations rather than their surface-level quality.
+G.2 Extended Limitations
+Context Length Constraints
+A practical limitation of LBMs is the context length restrictions of current LLMs.
+Even though modern models can handle thousands of tokens, comprehensive student histories in real educational settings can easily exceed these limits.
+As student trajectories grow more complex and include more information, the context required for the encoder might go beyond what current LLMs are capable of.
+A natural solution could be an iterative encoding process, where the encoder iterates over the text bottleneck while going over the input data by windows.
+Computational Cost and Resource Requirements
+Since LBMs involve two LLMs as the encoder and decoder, they are typically more computationally intensive to run than traditional KT methods.
+Each inference call with LBMs requires two LLM calls (encoding and decoding) with extensive context windows.
+Training LBMs with GRPO is particularly costly, as it requires iteratively generating and evaluating multiple candidate summaries per training example.
+This cost structure may limit the practical deployment of LBMs to high-stakes educational contexts where the interpretability benefits justify the computational expense.
+Dependence on Textual Question Content
+Many existing CD/KT datasets provide only question identifiers rather than full question text.
+This limits the direct applicability of LBMs, as they require the questions content in order to generate meaningful summaries about student knowledge.
+While question content is increasingly available in modern educational platforms, this dependency creates a barrier to applying LBMs to historical datasets or systems that rely primarily on item response theory frameworks.
+Constant Knowledge State Assumption
+The constant knowledge state assumption underlying the inverse problem formulation, while reasonable for short diagnostic sessions, does not cover longer time horizons.
+Real learning involves continuous knowledge acquisition, forgetting, and misconception evolution that our current framework cannot capture.
+This limitation restricts current LBMs to diagnostic rather than contexts of assessment throughout the learning process.
+The assumption also fails to account for contextual factors (fatigue, motivation, external stressors) that can significantly impact student performance within even short sessions.
+G.3 Extended Future Work
+Iterative Encoding Strategies
+A natural extension to address context length limitations involves iteratively applying the encoder while chunking input data.
+At each step, the encoder would process a portion of the interaction history $X$ along with one or more previously generated bottleneck summaries, creating an updated summary that incorporates new information from the latest chunk.
+This approach could maintain both detailed recent context and compressed historical patterns, enabling LBMs to handle arbitrarily long student trajectories through sequential refinement of knowledge state representations.
+Active Learning and Question Selection
+Building on iterative encoding, active learning strategies could provide encoders with the most informative input questions.
+Rather than processing all available interactions, the system could strategically select questions that maximize information gain about uncertain aspects of student knowledge.
+This could leverage uncertainty quantification methods or use the encoder LLM itself to recommend the most diagnostically valuable questions.
+Such active sensing would improve both efficiency and diagnostic power by focusing computational resources on the most insightful student responses.
+Dynamic Knowledge State Modeling
+Extending LBMs to handle evolving knowledge states represents a significant next step necessary to account for progressive learning and forgetting effects.
+A natural relaxation of the constant knowledge state assumption involves "piecewise-constant" knowledge states that can evolve between question sessions but remain static within sessions.
+This extension poses exciting challenges in developing training objectives that balance within-session consistency with between-session learning dynamics, potentially requiring new approaches to temporal modeling in natural language representations.
+Multimodal Input Integration
+Question-answer pairs represent a relatively narrow information source for inferring student knowledge states.
+Richer data sources such as student-tutor interactions, self-reported explanations for answers, timing patterns, or hint usage could provide deeper insights into student understanding.
+While LBMs naturally extend to text-based inputs, future work should investigate optimal strategies for combining these varied data sources and evaluate their relative contributions to knowledge state inference accuracy and interpretability.
+Pedagogical Alignment
+Recent work like LearnLM (Team et al., 2024) demonstrates the potential for making LLMs more pedagogically aligned.
+Incorporating similar pedagogical principles into LBM components could help encoders better interpret educational interactions and generate summaries that align with expert teaching practices.
+This might involve training on educator-annotated examples, incorporating educational taxonomies into summary structure, or using reward functions that emphasize pedagogically relevant aspects of student knowledge states.
+Such alignment could bridge the gap between computational convenience and educational validity.